### v1

# Infinite-Width Limit of a Single Attention Layer: Analysis via Tensor Programs

Abstract.
In modern theoretical analyses of neural networks, the infinite-width limit is often invoked to justify Gaussian approximations of neuron preactivations (e.g., via neural network Gaussian processes or Tensor Programs). However, these Gaussian-based asymptotic theories have so far been unable to capture the behavior of attention layers, except under special regimes such as infinitely many heads or tailored scaling schemes. In this paper, leveraging the Tensor Programs framework, we rigorously identify the infinite-width limit distribution of variables within a single attention layer under realistic architectural dimensionality and standard $1/\sqrt{n}$-scaling with $n$ dimensionality. We derive the exact form of this limit law without resorting to infinite-head approximations or tailored scalings, demonstrating that it departs fundamentally from Gaussianity. This limiting distribution exhibits non-Gaussianity from a hierarchical structure, being Gaussian conditional on the random similarity scores. Numerical experiments validate our theoretical predictions, confirming the effectiveness of our theory at finite width and accurate description of finite-head attentions. Beyond characterizing a standalone attention layer, our findings lay the groundwork for developing a unified theory of deep Transformer architectures in the infinite-width regime.
1. Introduction
A useful approach to understanding the complex probabilistic behavior of neural networks is through the study of parameter distributions in the infinite-width limit. Notable examples include the neural network Gaussian process (NNGP) LBN+ (18); HBSDN (20), which approximates the limit of stochastic parameter distributions with Gaussian processes; the neural tangent kernel (NTK) JGH (18), which represents the model near the initial value with kernel functions; mean field theory, which describes the update of parameter distributions MMN (18); and the Tensor Program, developed by Yan19a ; Yan20a ; YL (21); Yan20b ; YH (21); YL (23); YHB+ (21); YYZH (24), which is a general probabilistic analytical framework that unifies the representation of the infinite-width limit for a wide range of neural architectures and multiple layers. These methods provide probabilistic models that closely approximate the complex phenomena of neural networks.
One challenge in the studies is to properly represent the attention mechanisms used in Transformers  VSP+ (17), which frequently appear in recent large-scale architectures DCLT (19); AAA+ (23); ZZL+ (23). Unlike ordinary multi-layer perceptrons, an attention layer has interactions between query and key variables by multiplication, which makes the infinite-width limit distribution considerably more complex as shown in HBSDN (20), for example. To avoid this difficulty, the existing studies have mainly limited themselves to two special settings to compute the parameter distribution of the limit: (i) Infinite-head regime: HBSDN (20) considers the NNGP for the attention mechanism with infinite heads, resulting in a Gaussian approximation. (ii) $1/n$-scaling regime: the Tensor Programs Yan19b approximate the multiplication by changing the scale of the input variables for attention layers from $1/\sqrt{n}$ to $1/n$, where $n$ is a dimension of variables. However, these simplifications compromise the expressiveness and structure of actual attention mechanisms (and consequently, of Transformers). Specifically, the limit distribution at the infinite-head is not an effective approximation of the actual attention mechanisms because it differs significantly from the finite-head case. Also, the $1/n$-scaling regime makes all similarity scores converge to zero in the infinite-width limit, making the model equivalent to not measuring similarity between key and query vectors. Therefore, the limit parameter distribution of the attention mechanism is still under development.
In this study, we investigate the infinite-width limit distribution of outputs of a single attention layer under the common scaling and number of heads. To achieve this, we apply the Tensor Programs framework and analyze a new class of variables defined by the multiplication of intermediate variables, and derive a corresponding limit distribution. This new class of variables allows us to represent the dot-product score by the multiplication of keys and queries in the attention mechanism.
We summarize our contributions as follows:
  • Non-Gaussian limiting distribution: We study a distribution of outputs of a single attention layer with the $1/\sqrt{n}$-scaling and finite heads, and demonstrate that in the infinite-width limit, the output distribution converges to a hierarchical Gaussian distribution, which is a type of non-Gaussian. Specifically, the limiting distribution is a Gaussian conditional on the random similarity score, and this score variable itself converges to a Gaussian.
  • Consistency with numerical experiments: Our experiments justify that our theoretical limit distribution accurately captures the non-Gaussian behaviors exhibited by attention mechanisms. Specifically, even when the width is finite, our theory proves sufficiently accurate, provided that the dimension is large enough.
  • Novel proof technique with dot-products: We develop a novel proof technique focused on analyzing the similarity score variables by the dot-products of an attention layer. More concretely, we first show a convergence of the score variables to their limiting distributions, and then prove a conditional weak convergence of the outputs of an attention layer. This analysis characterizes the output distribution of an attention layer by incorporating the intrinsic randomness from the dot-product.
1.1. Related Works
The concentration of measures in neural networks plays a fundamental role in both theoretical analysis and practical applications of machine learning. Initially, studies in this area aimed to characterize the feedforward signal propagation in wide neural networks with random weights, inspired by statistical mechanics Ama (77); SCS (88). The outputs of such random neural networks converge to a Gaussian process, and the computation of signal propagation reduces to the composition of kernel functions of Gaussian processes utilized in machine learning Nea (96); Wil (96); DFS (16); LBN+ (18); dGMHR+ (18). They are often referred to as the NNGP. Since the NNGP kernel captures intrinsic inductive biases of architectures, its prediction performance correlates well with trained networks across different architectures LBN+ (18). Moreover, one can interpret NNGP as a network with random initialization of optimization, which naturally leads to quantitative insight into desirable weight scales to avoid exploding/vanishing signal and gradient problems PLR+ (16); SGGSD (17).
Depending on the network architecture, various types of kernels can be obtained, not only for fully-connected neural networks, but also convolutional neural networks (CNNs) NXL+ (19), skip connections YS (17), naive or gated recurrent neural networks (RNNs) CPS (18); Yan19b , and more YPR+ (19); GHLG (23). While some classical works assumed random Gaussian weights generated in an i.i.d. manner, recent research has shown that the same NNGP can be derived even in networks with non-Gaussian GY (22) or weakly correlated weights SNT (24). The NNGP can also be obtained for a network with a narrow bottleneck layer sandwiched between wide layers APH (20).
Two pioneering studies have investigated the NNGP of self-attention layers Yan19b ; HBSDN (20). Initially, Yan19b pointed out that an unconventional scaling factor of $1/n$ in the softmax function enables straightforward NNGP evaluation. Greg Yang has introduced the theoretical framework of Tensor Programs Yan19a ; Yan20a ; Yan20b , systematically composing kernel functions, including NNGPs, for modern neural networks. Transformers with the $1/n$-scaling fall within the scope of the applicability of Tensor Programs. However, the $1/n$-scaling is rarely used in theoretical or practical contexts, leaving the more realistic $1/\sqrt{n}$-scaling unresolved. To attack this problem, HBSDN (20) analyzed self-attention by varying the number of heads. They numerically demonstrated non-Gaussian behavior emerging in the single-head case due to stochasticity and correlations within the attention matrix, even at infinite embedding dimensions. They further showed that if we take the infinite limit of the number of heads, Gaussian behavior emerges in the self-attention output and defines an NNGP kernel termed infinite attention. Although this infinite attention empirically improved performance on certain NNGP regression benchmarks, its suitability as a theoretical foundation for realistic self-attention remains uncertain. This is because practical attention implementations typically use only 1-128 heads EXW+ (24), far fewer than the embedding dimension. In contrast to prior studies, our framework explicitly addresses scenarios with a finite number of heads and the resulting non-Gaussian effects.
2. Preliminary
2.1. Notation and Setup of Neural Networks
We define the notation for a standard neural network and its usage. In what follows, we denote by $n$ the dimensionality corresponding to the network’s width. Although we can vary $n$ across different layers or architectures, for simplicity we here treat every layer as having the same width $n$.
Neural network
We define notation for standard neural networks, excluding the attention mechanism. In particular, we adopt notations inspired by the framework of Tensor Programs Yan19a ; Yan19b , which allow us to describe a broad class of neural network architectures.
We describe an architecture of feed-forward neural networks as a finite set of $\mathbb{R}^{n}$-valued random vectors $h^{1},\dots,h^{J}$, which is inductively generated by the following rule. We fix a nonempty subset $\mathcal{V}_{0}\subset\{h^{1},\dots,h^{J}\}$, called the set of initial vectors (input layer). For each index $k$ with $h^{k}\notin\mathcal{V}_{0}$, the vector $h^{k}$ is generated either by matrix multiplication (MatMul) or by a coordinatewise nonlinearity (Nonlin). In the MatMul rule, given a weight matrix $W\in\mathbb{R}^{n\times n}$ and some $j\neq k$, one sets $h^{k}=Wh^{j}$. In the Nonlin rule, given $k\notin\{j_{1},\dots,j_{m}\}\subset[J]$ and a function $\phi:\mathbb{R}^{m}\to\mathbb{R}$, one sets
$h^{k}=\phi(h^{j_{1}},\dots,h^{j_{m}}),\quad h^{k}_{\alpha}=\phi(h^{j_{1}}_{\alpha},\dots,h^{j_{m}}_{\alpha})\quad(\alpha\in[n]).$
This type of Tensor Program is called Netsor, and it covers a broad range of neural network architectures, including a perceptron layer, a convolutional layer, a recurrent layer, and many others Yan19a ; Yan19b .
We present the perceptron layer as a specific neural network represented by Netsor. Suppose $x^{\ell-1}\in\mathbb{R}^{n}$ is the input of the $\ell$-th layer. It generates pre-activation variable $z^{\ell}\in\mathbb{R}^{n}$ and the input of the next layer $x^{\ell}\in\mathbb{R}^{n}$ by
$z^{\ell}=W^{\ell}x^{\ell-1},\quad x^{\ell}=\phi(z^{\ell}),$ (1)
where $W^{\ell}\in\mathbb{R}^{n\times n}$ is a weight matrix and $\phi:\mathbb{R}\to\mathbb{R}$ is a (potentially nonlinear) coordinatewise activation function. The vectors $z^{\ell}$ and $x^{\ell}$ are generated by MatMul and Nonlin, respectively.
Multi-head attention
We define the attention layer. Let $s$ and $H$ denote the spatial dimension and the number of heads, respectively. Suppose $W^{Q,a},W^{K,a},W^{V,a},W^{O,a}\in\mathbb{R}^{n\times n}$ are weight matrices for head $a\in[H]$. With an input sequence of $s$ random vectors $x^{1},\dots,x^{s}\in\mathbb{R}^{n}$, we define $X\in\mathbb{R}^{s\times n}$ as a matrix whose $i$-th row is $x^{i}$, i.e., $X^{\top}=[x^{1}\ \cdots\ x^{s}]$. For each head $a\in[H]$, define $Q^{(a)}=X(W^{Q,a})^{\top},K^{(a)}=X(W^{K,a})^{\top},V^{(a)}=X(W^{V,a})^{\top}$, so that $Q^{(a)},K^{(a)},V^{(a)}$ are $\mathbb{R}^{s\times n}$ matrices. The scaled dot-product score matrix is given by
$G^{(a)}=\frac{1}{\sqrt{n}}Q^{(a)}(K^{(a)})^{\top}=(p_{i,j}^{(a)})_{i,j\in[s]}\in\mathbb{R}^{s\times s},\quad p_{i,j}^{(a)}=\frac{1}{\sqrt{n}}(W^{Q,a}x^{i})^{\top}(W^{K,a}x^{j}).$ (2)
Applying the row-wise SoftMax to the score matrix, we define
$A^{(a)}=\mathrm{SoftMax}(G^{(a)})\in\mathbb{R}^{s\times s},\quad A^{(a)}_{i,j}=\mathrm{SoftMax}_{j}(p^{(a)}_{i,1},\dots,p^{(a)}_{i,s}).$ (3)
We then weight the values to form each head’s output $\mathrm{Head}^{(a)}=A^{(a)}V^{(a)}\in\mathbb{R}^{s\times n}$. Finally, we sum across all $H$ heads to recover the full attention output $\mathrm{MultiHead}=\sum_{a=1}^{H}\mathrm{Head}^{(a)}(W^{O,a})^{\top}\in\mathbb{R}^{s\times n}$, where each row $\mathrm{MultiHead}_{i\cdot}$ is given by
$(\mathrm{MultiHead}_{i\cdot})^{\top}=\sum_{a=1}^{H}\sum_{j=1}^{s}W^{O,a}W^{V,a}x^{i}\mathrm{SoftMax}_{j}(p^{(a)}_{i,1},\dots,p^{(a)}_{i,s})\in\mathbb{R}^{n}\quad(i\in[s]).$ (4)
Weight initialization
We set the weight matrices such as $W^{\ell},W^{Q,a},W^{K,a},W^{V,a}$, and $W^{O,a}$, and initial vectors $h\in\mathcal{V}_{0}$ as follows:
  1. (i)
    Each weight matrix $W$ is independent. Each $(\alpha,\beta)$ element of $W$ is sampled i.i.d. from $W_{\alpha\beta}\sim N(0,\sigma_{W}^{2}/n)$, where $\sigma_{W}>0$ is a constant that may depend on $W$.
  2. (ii)
    Let $Z^{\mathcal{V}_{0}}=\{Z^{h}:h\in\mathcal{V}_{0}\}\in\mathbb{R}^{|\mathcal{V}_{0}|}$ be a multivariate normal distribution. For each $\alpha\in[n]$, the collection of $\alpha$-th components of all initial vectors in $\mathcal{V}_{0}$, denoted by $\{h_{\alpha}:h\in\mathcal{V}_{0}\}$, is sampled i.i.d. from $Z^{\mathcal{V}_{0}}$.
This setup of weight matrices is common in practice (see GB (10) for summary) and also identical to that in the existing tensor networks Yan19a ; Yan19b ; Yan20b .
2.2. Distribution of Attention Mechanism in Previous Setup
We discuss the existing challenges on characterizing the output distribution of the attention layer, which exhibits non-Gaussian behavior. To derive this distribution, two regimes are typically made: the choice of scaling in the similarity computation, and the number of heads.
$1/n$-scaling regime
Yan19b studies an attention layer whose scaling term $1/\sqrt{n}$ in the scaled dot-product of Eq. (2) is replaced by $1/n$. In this regime, the Tensor Programs framework can show that the outputs of the attention layer converge to a Gaussian distribution: as $n\to\infty$, it holds that
$\textstyle\mathrm{MultiHead}_{i\alpha}\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(0,\kappa^{2}),\quad(\alpha\in[n]),$ (5)
with some $\kappa>0$ (see Appendix A and Theorem E.8 in Yan19b for details). Intuitively, in the $1/n$-scaling regime, the dot-product score $p_{i,j}^{(a)}$ converges to 00 for every pair $(i,j)\in[s]^{2}$ and head $a\in[H]$, which simplifies the non-Gaussian behavior of the attention outputs.
Infinite-heads regime
HBSDN (20) considers the regime in which many dimensionality diverge to infinite, then shows the convergence of the attention outputs to Gaussian. Specifically, it is shown that as $n,H\to\infty$, it holds that
$\textstyle\mathrm{MultiHead}_{i\alpha}\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(0,(\kappa^{\prime})^{2}),\quad(\alpha\in[n]),$ (6)
with some $\kappa^{\prime}>0$ (see Theorem 3 in HBSDN (20)). The variance $(\kappa^{\prime})^{2}$ is described by a covariance of the nonlinearly transformed dot-product score $p_{i,j}^{(a)}$. In the result, by letting the number of heads $H$ grow to infinite, the complex non-Gaussian effects from the dot-product score are smoothed out, yielding a convenient Gaussian distribution.
While these approximations are analytically appealing, they have notable limitations. In practice, Transformers typically use $1/\sqrt{n}$-scaling (see Equation (1) in VSP+ (17)) and a finite number of heads, resulting in the frequently observed non-Gaussian behaviors of attention layers. Although the above regimes simplify the behavior into a Gaussian form for tractability, a more precise theory is needed to capture the true behavior of attention layers in practice.
3. Main Theorem
3.1. Limiting Distribution
We introduce the limiting distribution for two types of quantities: variables generated within the Netsor program, and the scalar dot-products that arise in attention mechanisms. The convergence of these variables to their limiting distributions is formally established in our main result, Theorem 1.
Definition 1 (Limiting Distribution).
(A) Limiting distribution for vectors in the Netsor program: For each vector $h\in\mathbb{R}^{n}$ in the Netsor program, there exist a corresponding random variable $Z^{h}$ as follows:
  1. (i)
    If $h$ is as an initial vector from $\mathcal{V}_{0}$, then $Z^{h}$ follows the distribution specified in Section 2.1.
  2. (ii)
    If $g^{1},\dots,g^{k}$ are generated by MatMul, then $(Z^{g^{1}},\dots,Z^{g^{k}})$ is a zero-mean Gaussian vector. Specifically, for $g^{i}=W^{i}h^{i}$ and $g^{j}=W^{j}h^{j}$, their covariance is given by
    $\mathrm{Cov}(Z^{g^{i}},Z^{g^{j}})=\begin{cases}0&(\text{ if }W^{i}\text{ and }W^{j}\text{ are different matrices}),\\ \sigma_{W}^{2}\mathbb{E}[Z^{h^{i}}Z^{h^{j}}]&(\text{ if }W^{i}\text{ and }W^{j}\text{ are the same matrix}).\end{cases}$
  3. (iii)
    If $h$ is generated by Nonlin, $h=\phi(h^{1},\dots,h^{k})$, then $Z^{h}$ is defined as $Z^{h}=\phi(Z^{h^{1}},\dots,Z^{h^{k}})$.
(B) Limiting distribution for scalar dot-products: For each scalar $p_{i}=(g^{i,1})^{\top}g^{i,2}/\sqrt{n}$, where $g^{i,1}$ and $g^{i,2}$ are outputs of MatMul within the Netsor program, we define $\mathring{p}_{i}$ as follows:
  1. (iv)
    $(\mathring{p}_{1},\dots,\mathring{p}_{r})$ is a zero-mean Gaussian vector. This vector is statistically independent of all $Z^{h}$ variables defined in (A). The covariance of $(\mathring{p}_{1},\dots,\mathring{p}_{r})$ is given by
    $\mathrm{Cov}(\mathring{p}_{i},\mathring{p}_{k})=\mathbb{E}[Z^{g^{i,1}}Z^{g^{i,2}}Z^{g^{k,1}}Z^{g^{k,2}}].$
In Definition 1, the formulations for the limiting random variables $Z^{h}$ associated with initial vectors, MatMul outputs, and Nonlin outputs are developed by the existing Tensor Programs framework (Yan19b, ). Specifically, the limiting distribution of initial vectors follows that given in Section 2.1; variables generated by MatMul converge to Gaussian distribution; and variables generated by Nonlin are nonlinear transformation of the corresponding limiting distributions.
The distinct aspect of our analysis lies in the treatment of the limiting distribution for the scalar dot-products $(\mathring{p}_{1},\dots,\mathring{p}_{r})$, which represent pre-softmax attention scores. While the marginal Gaussian distribution of these scores is consistent with that in the infinite-head limit (see Theorem 3 in HBSDN (20)), our framework explicitly defines them as a Gaussian vector that is independent of the limiting distributions $Z^{h}$ within the Netsor program. This independence facilitates the computation of the limiting distribution of the attention outputs, as it allows us to treat the attention scores separately from the other variables in the Netsor program.
3.2. Statement
In this section, we show the convergence of the distribution of variables in the presence of an attention layer as the main result. Here, we refer to a function as pseudo-Lipschitz if it is pseudo-Lipschitz of order $d$ for some $d\in[2,\infty)$. See Definition 2 for the definition of pseudo-Lipschitz functions.
As a preparation, we first introduce the results of the convergence of the Netsor program without an attention layer. The following theorem is a slight simplification of Theorem 5.4 in Yan19b . [^1] Yan19b considers controlled functions instead of pseudo-Lipschitz functions, the former of which is a generalization of pseudo-Lipschitz functions.
Fact 1 (Netsor Master Theorem Yan19b ).
Consider a Netsor program. Suppose all initial vectors and weight matrices are sampled as explained in Section 2.1, and all nonlinearities used in Nonlin are pseudo-Lipschitz. For any positive integer $k$, let $h^{1},\dots,h^{k}$ be any vectors in the Netsor program. Then, for any pseudo-Lipschitz function $\psi:\mathbb{R}^{k}\to\mathbb{R}$, we have
$\frac{1}{n}\sum_{\alpha=1}^{n}\psi(h^{1},\dots,h^{k})\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\mathbb{E}[\psi(Z^{h^{1}},\dots,Z^{h^{k}})]$
as $n\to\infty$. Here, $Z^{h^{1}},\dots,Z^{h^{k}}$ are defined in Definition 1, (i)–(iii).
Building on the Netsor master theorem, we now present our main result. Before that, we introduce some assumptions specific to the attention layer.
Assumption 1.
Let $r$ and $m$ be positive integers that satisfy $m\geq 2r$. Suppose $g^{1},\dots,g^{m}\in\mathbb{R}^{n}$ are vectors in the Netsor program generated by MatMul. We assume that a subset $\{g^{i,j}:i\in[r],\ j\in[2]\}\subset\{g^{1},\dots,g^{m}\}$ can be expressed as, without loss of generality, [^2] The Tensor Programs framework generates vectors inductively, precluding circular dependencies (see Section 2.1). Consequently, for each $x^{i,j}=\phi^{i,j}(g^{1},\dots,g^{m})$, any arguments from $\{g^{1},\dots,g^{m}\}$ that effectively contribute to the output $x^{i,j}$ (i.e., those upon which the function $\phi^{i,j}$ actually depends) must be defined prior to $g^{i,j}=W^{i,j}x^{i,j}$ in this inductive sequence.
$g^{i,j}=W^{i,j}x^{i,j},\quad x^{i,j}=\phi^{i,j}(g^{1},\dots,g^{m})\quad(i\in[r],\ j\in[2]),$
where each $\phi^{i,j}$ is a bounded and pseudo-Lipschitz function. The weight matrices $W^{i,j}\in\mathbb{R}^{n\times n}$ satisfy two conditions:
  1. (a)
    $\{W^{i,j}:i\in[r],\ j\in[2]\}$ is specific to the generation of $\{g^{i,j}:i\in[r],\ j\in[2]\}$ and is not used for any $g\in\{g^{1},\dots,g^{m}\}\setminus\{g^{i,j}:i\in[r],\ j\in[2]\}$.
  2. (b)
    It is permissible for $W^{i,j}$ to be the same matrix as $W^{i^{\prime},j^{\prime}}$ unless $i=i^{\prime},j\neq j^{\prime}$ (e.g., $W^{1,1}$ could be the same matrix as $W^{2,1}$).
Finally, we define the scalar dot-products $\{p_{i}:i\in[r]\}$ by
$p_{i}=\frac{1}{\sqrt{n}}(g^{i,1})^{\top}g^{i,2}\quad(i\in[r]).$
Theorem 1.
Consider a Netsor program, and suppose all nonlinearities used in Nonlin are pseudo-Lipschitz. We adopt the settings and notations from Assumption 1, which defines vectors $g^{1},\dots,g^{m}$ and scalar dot-products $p_{1},\dots,p_{r}$. Further, suppose all initial vectors and weight matrices are sampled as explained in Section 2.1. Now, let $h^{1},\dots,h^{k}\in\mathbb{R}^{n}$ be vectors whose elements are given by
$h_{\alpha}^{j}=\varphi^{j}(g_{\alpha}^{1},\dots,g_{\alpha}^{m},p_{1},\dots,p_{r})\quad(\alpha\in[n],\ j\in[k]),$
where each $\varphi^{j}$ is a pseudo-Lipschitz function. Then, for any bounded and pseudo-Lipschitz function $\psi:\mathbb{R}^{k}\to\mathbb{R}$, we have
$\frac{1}{n}\sum_{\alpha=1}^{n}\psi(h_{\alpha}^{1},\dots,h_{\alpha}^{k})\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathbb{E}[\psi(Z^{h^{1}},\dots,Z^{h^{k}})\mid\mathring{p}_{1},\dots,\mathring{p}_{r}]$
as $n\to\infty$. Here, $Z^{h^{j}}$ is given by $Z^{h^{j}}=\varphi^{j}(Z^{g^{1}},\dots,Z^{g^{m}},\mathring{p}_{1},\dots,\mathring{p}_{r})$ for $j\in[k]$, and $(Z^{g^{1}},\dots,Z^{g^{m}},\mathring{p}_{1},\dots,\mathring{p}_{r})$ is defined in Definition 1.
It may be noted that in the statement of Theorem 1, the boundedness of $\phi^{i,j}$ as in Assumption 1 is not essential, and we can drop this condition. However, we keep it for simplicity of the proof. Also, note that in this statement, $\mathring{p}_{1},\dots,\mathring{p}_{r}$ are random variables with positive variance, and the convergence in distribution ${\to}^{d}$ described in the statement refers to convergence to the distribution of them. Unlike the conventional master theorem derived from Tensor Programs Yan19a ; Yan19b , this theorem simultaneously characterizes the randomness of the ordinary variables like $g^{1},\dots,g^{m}$, along with the effects of finite-dimensional random variables $\mathring{p}_{1},\dots,\mathring{p}_{r}$.
Theorem 1 implies that, due to the randomness of $(\mathring{p}_{1},\dots,\mathring{p}_{r})$, the overall outputs resembles a hierarchical distribution, where the variance of one limiting distribution may depend on a realization of another limiting distribution $(\mathring{p}_{1},\dots,\mathring{p}_{r})$. This result differs from existing results by the Tensor Programs where the variance depends on moments of the other distributions. This result leads to a non-Gaussian limiting distribution in the attention case, whose details will be provided in Example 1.
The following corollary is an analogue of Theorem A.5 in Yan20b .
Corollary 2 (Coordinatewise Convergence).
Assume the same premise as in Theorem 1. Then, for all $\alpha\geq 1$, we have
$(h_{\alpha}^{1},\dots,h_{\alpha}^{k})\stackrel{{\scriptstyle d}}{{\longrightarrow}}(Z^{h^{1}},\dots,Z^{h^{k}}).$
We present specific multi-head attention results as an example of the application of Corollary 2. This example implies that the output of the attention mechanisms is described by a non-Gaussian distribution. Specifically, when $y^{i}$ is an output of a multi-head attention layer as $\mathrm{MultiHead}_{i\cdot}$ defined in Eq. (4), $(Z^{y^{1}},\dots,Z^{y^{s}})$ follows a hierarchical Gaussian distribution whose conditioning $(\mathring{p}_{1},\dots,\mathring{p}_{r})=\{\mathring{p}_{i,j}^{(a)}:i,j\in[s],\ a\in[H]\}$ itself is a random variable, causing $(Z^{y^{1}},\dots,Z^{y^{s}})$ to converge to a heavy-tailed non-Gaussian distribution overall. We provide the details as follows:
Example 1 (Multi-Head Attention).
We consider the multi-head attention in Eq. (4). Recall that $s$ is the spatial dimension and $H$ is the number of heads. We sample new weight matrices $W^{Q,a},W^{K,a},W^{V,a},W^{O,a}\in\mathbb{R}^{n\times n}$ for $a\in[H]$. Let $x^{1},\dots,x^{s}\in\mathbb{R}^{n}$ be vectors within the Netsor program. We assume these input vectors are generated by Nonlin, where the nonlinearity is bounded and pseudo-Lipschitz. For $j\in[s],\ a\in[H]$, we define the value vectors $v^{a,j}=W^{V,a}x^{j}\in\mathbb{R}^{n}\quad(j\in[s],\ a\in[H])$ and its further transform $\tilde{v}^{a,j}=W^{O,a}v^{a,j}\in\mathbb{R}^{n}$. Finally, with the score $p_{i,j}^{(a)}$ in Eq. (2), we rewrite an element of the multi-head attention Eq. (4) as
$y_{\alpha}^{i}$ $=\sum_{a=1}^{H}\sum_{j=1}^{s}\mathrm{SoftMax}_{j}(p_{i,1}^{(a)},\dots,p_{i,s}^{(a)})\tilde{v}_{\alpha}^{a,j}$
$=:\varphi^{i}\left(\{\tilde{v}_{\alpha}^{a,j}:j\in[s],\ a\in[H]\},\ \{p_{i,j}^{(a)}:i,j\in[s],\ a\in[H]\}\right),$
by introducing functions $\varphi^{i}\ (i\in[s])$, each of which is a pseudo-Lipschitz function. [^3] See Proposition 6 and Lemma 7. Note that $\mathrm{SoftMax}$ is Lipschitz so that it is also pseudo-Lipschitz (Fact 4). Then, by Corollary 2, we have
$(y_{\alpha}^{1},\dots,y_{\alpha}^{s})\stackrel{{\scriptstyle d}}{{\longrightarrow}}(Z^{y^{1}},\dots,Z^{y^{s}})\quad(n\to\infty),$
where $Z^{y^{i}}$ is defined by $Z^{y^{i}}=\sum_{a=1}^{H}\sum_{j=1}^{s}\mathrm{SoftMax}_{j}(\mathring{p}_{i,1}^{(a)},\dots,\mathring{p}_{i,s}^{(a)})Z^{\tilde{v}^{a,j}}$. By Definition 1, $\{Z^{\tilde{v}^{a,j}}:j\in[s],\ a\in[H]\}$ and $\{\mathring{p}_{i,j}^{(a)}:i,j\in[s],\ a\in[H]\}$ satisfy the following.
  • $\{Z^{\tilde{v}^{a,j}}:j\in[s],\ a\in[H]\}$ is independent of $\{\mathring{p}_{i,j}^{(a)}:i,j\in[s],\ a\in[H]\}$.
  • $\{Z^{\tilde{v}^{a,j}}:j\in[s],\ a\in[H]\}$ is jointly Gaussian with zero mean. For $a,a^{\prime}\in[H]$ and $j,j^{\prime}\in[s]$, the covariance is given by
    $\mathrm{Cov}(Z^{\tilde{v}^{a,j}},Z^{\tilde{v}^{a^{\prime},j^{\prime}}})=\begin{cases}0&(a\neq a^{\prime}),\\ \sigma_{W^{O,a}}^{2}\sigma_{W^{V,a}}^{2}\mathbb{E}[Z^{x^{j}}Z^{x^{j^{\prime}}}]&(a=a^{\prime}).\end{cases}$
  • $\{\mathring{p}_{i,j}^{(a)}:i,j\in[s],\ a\in[H]\}$ is jointly Gaussian with zero mean. For $a,a^{\prime}\in[H]$ and $i,i^{\prime},j,j^{\prime}\in[s]$, the covariance is given by
    $\mathrm{Cov}(\mathring{p}_{i,j}^{(a)},\mathring{p}_{i^{\prime},j^{\prime}}^{(a^{\prime})})=\begin{cases}0&(a\neq a^{\prime}),\\ \sigma_{W^{Q,a}}^{2}\sigma_{W^{K,a}}^{2}\mathbb{E}[Z^{x^{i}}Z^{x^{i^{\prime}}}]\mathbb{E}[Z^{x^{j}}Z^{x^{j^{\prime}}}]&(a=a^{\prime}).\end{cases}$
Note that $\{Z^{x^{j}}:j\in[s]\}$ can be computed by Definition 1.
4. Proof Sketch of Theorem 1
We sketch the proof of Theorem 1. First, observe that it is equivalent to the following statement:
Theorem 3.
Assume the same premise as in Theorem 1. Then, for any bounded and pseudo-Lipschitz function $\psi$, we have
$\frac{1}{n}\sum_{\alpha=1}^{n}\psi(g_{\alpha}^{1},\dots,g_{\alpha}^{m},p_{1},\dots,p_{r})\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathbb{E}[\psi(Z^{g^{1}},\dots,Z^{g^{m}},\mathring{p}_{1},\dots,\mathring{p}_{r})\mid\mathring{p}_{1},\dots,\mathring{p}_{r}]$
as $n\to\infty$, where $(Z^{g^{1}},\dots,Z^{g^{m}},\mathring{p}_{1},\dots,\mathring{p}_{r})$ is defined in Definition 1.
Clearly, Theorem 1 implies Theorem 3. Conversely, applying Proposition 6 establishes the opposite direction.
In what follows, we outline the proof of Theorem 3. Hereafter, $\bm{x}_{1:k}$ denotes the vector $(x_{1},\dots,x_{k})$, and likewise, $\bm{x}_{\alpha}^{1:k}$ denotes the vector $(x_{\alpha}^{1},\dots,x_{\alpha}^{k})$. Fix a bounded and pseudo-Lipschitz function $\psi$. By Lemma 5, it suffices to show that, for any bounded and Lipschitz function $f$,
$\left|\mathbb{E}f\left(\frac{1}{n}\sum_{\alpha=1}^{n}\psi(\bm{g}_{\alpha}^{1:m},\bm{p}_{1:r})\right)-\mathbb{E}f\left(\mathbb{E}[\psi(Z^{\bm{g}^{1:m}},\mathring{\bm{p}}_{1:r})\mid\mathring{\bm{p}}_{1:r}]\right)\right|\to 0$
holds. Note that by Fact 1, we have
$\frac{1}{n}\sum_{\alpha=1}^{n}\tilde{\psi}(\bm{g}_{\alpha}^{1:m})\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\mathbb{E}[\tilde{\psi}(Z^{\bm{g}^{1:m}})]$ (7)
for any pseudo-Lipschitz function $\tilde{\psi}$. Using this result, we first show that $\bm{p}_{1:r}$ converges in distribution to $\mathring{\bm{p}}_{1:r}$, which is a Gaussian vector defined by Definition 1 (Appendix C.1.1). Then, we consider
$\left|\mathbb{E}f\left(\frac{1}{n}\sum_{\alpha=1}^{n}\psi(\bm{g}_{\alpha}^{1:m},\bm{p}_{1:r})\right)-\mathbb{E}f\left(\mathbb{E}[\psi(Z^{\bm{g}^{1:m}},\mathring{\bm{p}}_{1:r})\mid\mathring{\bm{p}}_{1:r}]\right)\right|\leq S_{1}+S_{2},$
where $S_{1}$ and $S_{2}$ are given by
$S_{1}$ $=\left|\mathbb{E}f\left(\frac{1}{n}\sum_{\alpha=1}^{n}\psi(\bm{g}_{\alpha}^{1:m},\bm{p}_{1:r})\right)-\mathbb{E}f\left(\mathbb{E}\left[\psi(Z^{\bm{g}^{1:m}},\bm{p}_{1:r})\mid\bm{p}_{1:r}\right]\right)\right|,$
$S_{2}$ $=\left|\mathbb{E}f\left(\mathbb{E}\left[\psi(Z^{\bm{g}^{1:m}},\bm{p}_{1:r})\mid\bm{p}_{1:r}\right]\right)-\mathbb{E}f\left(\mathbb{E}[\psi(Z^{\bm{g}^{1:m}},\mathring{\bm{p}}_{1:r})\mid\mathring{\bm{p}}_{1:r}]\right)\right|.$
We separately show $S_{1}\to 0$ (Appendix C.1.2) and $S_{2}\to 0$ (Appendix C.1.3).
5. Simulation and Discussion
We perform simulations to validate the infinite-width limit distributions derived in Example 1. For simplicity, the spatial dimension is fixed at $s=4$ in all experiments. More details on the simulation setting can be found in Appendix A. All simulation codes are available at https://github.com/manasakai/infinite-width-attention.
5.1. Effect of Finite Width
(a) Distribution of $y_{1}^{1}$ and $Z^{y^{1}}$.
(b) KL divergence with error bars.
Figure 1. Comparison of the distribution of the attention output $y_{1}^{1}$ and its infinite-width limit $Z^{y^{1}}$ in Example 1. (A) Kernel density estimates of the empirical distribution (via Monte Carlo sampling) of $y_{1}^{1}$ for widths $n\in\{16,64,256,1024\}$ (dashed lines) alongside that of $Z^{y^{1}}$ (solid line), showing the convergence of the finite-width distribution to its limit. (B) Average of the log-KL divergence $\log\mathrm{KL}(\mathrm{Dist}(y_{1}^{1})\|\mathrm{Dist}(Z^{y^{1}}))$ over 10 independent trials, plotted against $\log_{4}(n)$ with error bars indicating one standard deviation, confirming a decreasing trend.
To assess how well the infinite-width theory of Theorem 1 aligns with finite-width multi-head attention behavior and to verify that discrepancies diminish as width grows, we perform 10 independent experiments for each width $n\in\{16,64,256,1024\}$. In each trial, we draw samples of the multi-head output $y_{1}^{1}$ with $H=2$ in Example 1, estimate its density via kernel density estimation, and compute the KL divergence to its theoretical limit $Z^{y^{1}}$, the density of which is also approximated via Monte Carlo sampling. Figure 1(A) plots the estimated densities of our first trial, showing that the density of $y_{1}^{1}$ converges rapidly to that of $Z^{y^{1}}$ as $n$ increases. Figure 1(B) quantifies this convergence by plotting the average log-KL divergence against $\log_{4}(n)$, with error bars showing one standard deviation across the 10 trials, demonstrating a consistent decay with growing width.
5.2. Scalings of the Dot-product Score and Finiteness of Heads
(a) Dot-product score under different scalings.
(b) Varying head count of the attention output.
Figure 2. Visualization of the dot-product score $p_{1,1}^{(1)}$ and attention output $y_{1}^{1}$, as defined in Example 1, comparing finite-width behavior to their infinite-width limits. (A) Histogram of the empirical distribution of $p_{1,1}^{(1)}$ for $n=256$ alongside the plot of its infinite-width limit distribution $\mathring{p}_{1,1}^{1}$, under two scaling schemes; $1/\sqrt{n}$ and $1/n$. The $1/n$-scaled score collapses to zero in the infinite-width limit, while the $1/\sqrt{n}$-scaled score retains a nondegenerate distribution. (B) Histogram of the empirical distribution of $y_{1}^{1}$ for $n=256$ alongside the plot of its infinite-width limit distribution $Z^{y^{1}}$, varying head counts $H\in\{1,256\}$. This demonstrates that our theoretical prediction remains accurate even when $H$ grows.
We next investigate two facets of finite-width behavior in Example 1, which are the effect of different scaling rules on the variability of the dot-product score and the impact of a finite number of heads on the attention output. For comparison, Yan19b assumes $1/n$-scaling, and the infinite-head result by HBSDN (20) applies only in the limit $n,H\to\infty$.
Figure 2(A) shows the histogram of the empirical distribution of the dot-product score $p_{1,1}^{(1)}$ at width $n=256$ under the two scaling schemes $1/\sqrt{n}$ and $1/n$, together with the infinite-width limit distribution $\mathring{p}_{1,1}^{1}$. Even at this moderate width, the $1/n$-scaled scores are tightly concentrated around zero, whereas the $1/\sqrt{n}$-scaled scores remain spread out in a nondegenerate fashion, confirming that only the latter preserves variability away from zero at large $n$.
Figure 2(B) reports histograms of the attention output $y_{1}^{1}$ for $n=256$ with head counts $H=1$ and $H=256$, overlaid with the infinite-width densities of $Z^{y^{1}}$, which are approximated via Monte Carlo sampling. The close agreement between finite-width histograms and their infinite-width limit curves, even when $H$ grows on the same order as $n$, demonstrates the robustness of our infinite-width approximation in the growing number of heads.
6. Conclusion
In this paper, we rigorously analyzed the infinite-width limit distribution of outputs from a single attention layer using the Tensor Programs framework. Specifically, we theoretically showed and empirically confirmed that the attention outputs converge to a non-Gaussian distribution under realistic conditions with finite heads and standard $1/\sqrt{n}$-scaling. Our results lay a theoretical foundation for handling attention mechanisms in practical settings and open the door for future extension to deep Transformer architectures. As a limitation of this study, extending these results to multi-layer structures remains challenging; particularly, a dynamic analysis on an attention layer, akin to analysis within the Tensor Programs framework, poses additional difficulties.
7. Acknowledgements
Mana Sakai was supported by RIKEN Junior Research Associate Program. Ryo Karakida was supported by JST FOREST (Grant No. JPMJFR226Q) and JSPS KAEKENHI (Grant No. 22H05116). Masaaki Imaizumi was supported by JSPS KAKENHI (Grant No. 24K02904), JST CREST (Grant No. JPMJCR21D2), and JST FOREST (Grant No. JPMJFR216I).
Appendix A Simulation Details and Supplementary Analysis
A.1. General Experimental Setup
Throughout all simulations presented in this paper, we consistently set the spatial dimension to $s=4$. The core experimental setup follows that described in Example 1. Specifically, let $x^{1},\dots,x^{s}$ be outputs of Nonlin. These are obtained by applying a clipping activation function $\psi$ to preceding vectors $h^{1},\dots,h^{s}\in\mathbb{R}^{n}$:
$x_{\alpha}^{i}=\psi(h_{\alpha}^{i})=-C1\{h_{\alpha}^{i}<-C\}+h_{\alpha}^{i}1\{-C\leq h_{\alpha}^{i}\leq C\}+C1\{h_{\alpha}^{i}>C\}\quad(\alpha\in[n],\ i\in[s]),$
where $1\{\cdot\}$ denotes the indicator function and $C$ is a positive constant. The vectors $h^{1},\dots,h^{s}$ are outputs of MatMul, defined by
$h^{i}=W^{i}h\quad(i\in[s]).$
Each element of the initial vector $h\in\mathbb{R}^{n}$ is sampled independently from a standard normal distribution.
For all weight matrices involved in the attention mechanism $W^{Q,a},W^{K,a},W^{V,a},W^{O,a}$ and the matrices $W^{i}$ generating $x^{i}$, we set
$\sigma_{W^{Q,a}}^{2}=\sigma_{W^{K,a}}^{2}=\sigma_{W^{V,a}}^{2}=\sigma_{W^{O,a}}^{2}=\sigma_{W^{i}}^{2}=1.$
The elements of these weight matrices are independently sampled from $\mathcal{N}(0,\sigma_{W}^{2}/n)$, where $\sigma_{W}^{2}$ is the respective variance (here, 1).
Under this setup, the input vectors $x^{j}$ to the attention layer are designed such that their infinite-width limits $Z^{x^{j}}$ are uncorrelated for $j\neq j^{\prime}$, i.e.,
$\mathbb{E}[Z^{x^{j}}Z^{x^{j^{\prime}}}]=0\quad(j,j^{\prime}\in[s],\ j\neq j^{\prime}).$
Furthermore, the infinite-width limit $Z^{h^{i}}$ is a standard normal random variable, leading to
$\mathbb{E}[(Z^{x^{i}})^{2}]=\mathbb{E}[(\psi(Z^{h^{i}}))^{2}]=2C^{2}(1-\Phi(C))-2C\phi(C)+2\Phi(C)-1,$
where $C$ is the constant used in the clipping activation, and $\Phi$ and $\phi$ are the cumulative distribution function and probability density function of the standard normal distribution, respectively. As $C\to\infty$, this value converges to $1$. In our experiments, we set $C=100$, and we approximate $\mathbb{E}[(Z^{x^{i}})^{2}]$ by $1$. Consequently, the covariances of the limiting variables for the vectors $\tilde{v}^{a,j}$ and dot-product scores $\mathring{p}_{i,j}^{(a)}$ simplify as described in Example 1,
$\mathrm{Cov}(Z^{\tilde{v}^{a,j}},Z^{\tilde{v}^{a^{\prime},j^{\prime}}})=\begin{cases}\mathbb{E}[(Z^{x^{j}})^{2}]\approx 1&(a=a^{\prime},\ j=j^{\prime}),\\ 0&(\text{otherwise}),\end{cases}$
and
$\mathrm{Cov}(\mathring{p}_{i,j}^{(a)},\mathring{p}_{i^{\prime},j^{\prime}}^{(a^{\prime})})=\begin{cases}\left(\mathbb{E}[(Z^{x^{i}})^{2}]\right)^{2}\approx 1&(a=a^{\prime},\ i=i^{\prime},\ j=j^{\prime}),\\ 0&(\text{otherwise}).\end{cases}$
To estimate the empirical distributions of finite-width attention outputs and their corresponding infinite-width limits, we employ Monte Carlo sampling. For each such estimation, $50,000$ samples are drawn. Kernel density estimation (KDE) is used to visualize these empirical distributions.
A.2. Analysis of Low-Rank Attention
A.2.1. Specific Setup for Low-Rank Attention
In practice, large-scale Transformers typically assume a specific embedding dimensionality for multi-head self-attention layers. For head counts $H$, the embedding dimension $n$ is set linearly as $n=Hn_{H}$, where $n_{H}$ denotes the head dimension and determines the sizes of weight matrices as $W^{Q,a},W^{K,a},W^{V,a}\in\mathbb{R}^{n_{H}\times n}$. Thus, the QK product becomes low-rank relative to the embedding dimension $n$, and the scaling factor is given by $1/\sqrt{n_{H}}$ as follows:
$p_{i,j}^{(a)}=\frac{1}{\sqrt{n_{H}}}(W^{Q,a}x^{i})^{\top}(W^{K,a}x^{j})\quad(i,j\in[s],\ a\in[H]).$
For example, the original Transformer architecture sets $H=8$ and $n_{H}=64$. Large-scale models often increase the number of heads to be on the order of the hidden embedding dimension EXW+ (24), as seen in GPT-3 (175B parameters), which sets $H=96$ and $n_{H}=128$.
Additionally, since the output from each head is also of dimension $n_{H}$ through the value matrix, an output weight $W^{O,a}\in\mathbb{R}^{n\times n_{H}}$ is applied to map it back to the $n$-dimensional input for the subsequent layer:
$y^{i}=\sum_{a=1}^{H}\sum_{j=1}^{s}\mathrm{SoftMax}_{j}(p_{i,1}^{(a)},\dots,p_{i,s}^{(a)})W^{O,a}W^{V,a}x^{j}\quad(i\in[s]).$
Note that, to ensure the dot-product scores and the attention outputs are of order 1, the weight matrices are randomly initialized with the following scales:
$W_{\alpha\beta}^{Q,a},W_{\alpha\beta}^{K,a},W_{\alpha\beta}^{V,a}\sim N(0,\sigma_{W}^{2}/n),\qquad W_{\alpha\beta}^{O,a}\sim N(0,\sigma_{W}^{2}/n_{H})\quad(\alpha,\beta\in[n]).$
A.2.2. Results and Discussion
Figure 3. Comparison of the distribution of the attention output $y_{1}^{1}$ under the low-rank setting and its infinite-width limit $Z^{y}$. Kernel density estimates of the empirical distribution (via Monte Carlo sampling) of $y_{1}^{1}$ for various widths $n$ and head counts $H$ (with $n_{H}=n/H=64$ is fixed, dashed lines) alongside that of $Z^{y}$ (solid lines). The plot shows convergence to the infinite-width limit as $n$ and $H$ increase proportionally.
Finally, we investigate the behavior of the attention output $y^{1}$ in the low-rank regime described above, where the number of heads $H$ increases proportionally with $n$. Fixing the head dimension $n_{H}=n/H=64$, Figure 3 presents kernel density estimates of the distribution of $y_{1}^{1}$ for $(n,H)\in\{(64,1),(256,4),(1024,16)\}$, alongside the distributions of their corresponding infinite-width limit $Z^{y}$, approximated via Monte Carlo sampling. In all cases, the finite-width estimates (dashed lines) closely track the infinite-width limits (solid lines).
Notably, even in these practically relevant settings employing low-rank attention (where head-specific projections are $n_{H}\times n$ or $n\times n_{H}$ rather than the $n\times n$ matrices primarily assumed in our formal derivations in Theorem 1), our infinite-width framework continues to provide an excellent approximation. This agreement suggests that the core principles of convergence captured by our theory extend robustly to common architectural variants like low-rank attention (with appropriate scaling considerations), underscoring the practical utility of our theoretical predictions under these structural assumptions common in modern attention models."
Appendix B Mathematical Tools
B.1. Basics
Lemma 4.
For $1\leq m\leq\infty$ and $a_{1},\dots,a_{k}\in\mathbb{R}$, we have
$\left|\sum_{i=1}^{k}a_{i}\right|^{m}\leq\left(\sum_{i=1}^{k}|a_{i}|\right)^{m}\leq k^{m-1}\sum_{i=1}^{k}|a_{i}|^{m}.$
Proof.
The first inequality is an application of the triangle inequality. The second inequality follows from Jensen’s inequality. Since $m\geq 1$, the function $x\mapsto x^{m}$ on $[0,\infty)$ is convex, and thus Jensen’s inequality implies
$\left(\sum_{i=1}^{k}|a_{i}|\right)^{m}=k^{m}\left(\frac{1}{k}\sum_{i=1}^{k}|a_{i}|\right)^{m}\leq k^{m}\frac{1}{k}\sum_{i=1}^{k}|a_{i}|^{m}=k^{m-1}\sum_{i=1}^{k}|a_{i}|^{m}$
as desired. ∎
Lemma 5 (Portmanteau lemma (Lemma 2.2 in Vaa (98))).
The following conditions are equivalent.
  1. (i)
    $X_{n}\stackrel{{\scriptstyle d}}{{\longrightarrow}}X$.
  2. (ii)
    $\mathbb{E}[f(X_{n})]\to\mathbb{E}[f(X)]$ for all bounded and continuous function $f$.
  3. (iii)
    $\mathbb{E}[f(X_{n})]\to\mathbb{E}[f(X)]$ for all bounded and Lipschitz function $f$.
  4. (iv)
    $\mathrm{P}(X_{n}\in B)\to\mathrm{P}(X\in B)$ for all Borel sets $B$ with $\mathrm{P}(X\in\delta B)=0$, where $\delta B$ denotes the boundary of $B$.
Fact 2.
Suppose $\{X_{n}\}_{n\in\mathbb{N}}$ is a sequence of integrable random variables that converges in probability to $X$. Then the following statements are equivalent.
  1. (i)
    The sequence $\{X_{n}\}_{n\in\mathbb{N}}$ is uniformly integrable.
  2. (ii)
    $\mathbb{E}(|X_{n}|)\to\mathbb{E}(|X|)<\infty$.
Remark 1.
If $X_{n}$ converges to 00 in probability, then by Fact 2, we have
$\mathbb{E}(|X_{n}|)=o(1)\quad\iff\quad\{X_{n}\}_{n\in\mathbb{N}}\text{ is uniformly integrable}.$
Moreover, since $|\mathbb{E}(X_{n})|\leq\mathbb{E}(|X_{n}|)$, it follows that $\mathbb{E}(X_{n})=o(1)$.
Fact 3.
Suppose there exists $\delta>1$ such that $\sup_{n}\mathbb{E}(|X_{n}|^{\delta})<\infty$. Then the sequence $\{X_{n}\}_{n\in\mathbb{N}}$ is uniformly integrable.
B.2. Pseudo-Lipschitz Functions
Definition 2 (Pseudo-Lipschitz functions (BM, 11)).
Let $d>1$ be a constant. A function $f:\mathbb{R}^{k}\to\mathbb{R}$ is pseudo-Lipschitz of order $d$ if there exists a constant $C>0$ such that, for all $x,y\in\mathbb{R}^{k}$,
$|f(x)-f(y)|\leq C\|x-y\|(1+\|x\|^{d-1}+\|y\|^{d-1})$
holds.
Fact 4.
The following statements hold.
  1. (i)
    A Lipschitz function is pseudo-Lipschitz of order $d$ for all $d>1$.
  2. (ii)
    A pseudo-Lipschitz function (of any given order) is continuous.
In this paper, we refer to a function as pseudo-Lipschitz if it is pseudo-Lipschitz of order $d$ for some $d\in[2,\infty)$.
Proposition 6.
Suppose $f:\mathbb{R}^{k}\to\mathbb{R}$ and $g_{i}:\mathbb{R}^{\ell}\to\mathbb{R}$ $(i\in[k])$ are pseudo-Lipschitz. Then the function $h:\mathbb{R}^{\ell}\to\mathbb{R}$ defined by $h(x)=f(g_{1}(x),\dots,g_{k}(x))$ is also pseudo-Lipschitz.
Proof.
Suppose $f$ is pseudo-Lipschitz of order $d_{0}+1$ and each $g_{i}$ is pseudo-Lipschitz of order $d_{i}+1$. Define a function $g:\mathbb{R}^{\ell}\to\mathbb{R}^{k}$ and a constant $d\geq 1$ by
$g(x)=(g_{1}(x),\dots,g_{k}(x)),\quad d=\max\{d_{1},\dots,d_{k}\}.$
Applying the pseudo-Lipschitz bounds for $f$ and each $g_{i}$ gives
$|h(x)-h(x^{\prime})|\lesssim\|g(x)-g(x^{\prime})\|\left(1+\|g(x)\|^{d_{0}}+\left\|g(x^{\prime})\right\|^{d_{0}}\right)$
and
$|g_{i}(x)-g_{i}(x^{\prime})|\lesssim\|x-x^{\prime}\|\left(1+\|x\|^{d_{i}}+\|x^{\prime}\|^{d_{i}}\right)\lesssim\|x-x^{\prime}\|\left(1+\|x\|^{d}+\|x^{\prime}\|^{d}\right).$
The last inequality implies
$\|g(x)-g(x^{\prime})\|=\left(\sum_{i=1}^{k}|g_{i}(x)-g_{i}(x^{\prime})|^{2}\right)^{1/2}\lesssim\|x-x^{\prime}\|\left(1+\|x\|^{d}+\|x^{\prime}\|^{d}\right).$
On the other hand, since
$\|g(x)\|^{2}=\sum_{i=1}^{k}|g_{i}(x)|^{2}\leq\left(\sum_{i=1}^{k}|g_{i}(x)|\right)^{2}$
holds in general, Lemma 4 implies
$\|g(x)\|^{d_{0}}\leq\left(\sum_{i=1}^{k}|g_{i}(x)|\right)^{d_{0}}\lesssim\sum_{i=1}^{k}|g_{i}(x)|^{d_{0}}.$
The pseudo-Lipschitz property of each $g_{i}$ yields
$|g_{i}(x)|\leq|g_{i}(0)|+|g_{i}(x)-g_{i}(0)|\lesssim 1+\|x\|(1+\|x\|^{d_{i}})\lesssim 1+\|x\|^{d_{i}+1}\lesssim 1+\|x\|^{d+1}.$
Hence, by Lemma 4, we have
$\|g(x)\|^{d_{0}}\lesssim(1+\|x\|^{d+1})^{d_{0}}\lesssim 1+\|x\|^{d_{0}(d+1)}.$
Combining these elements gives
$|h(x)-h(x^{\prime})|$ $\lesssim\|x-x^{\prime}\|\left(1+\|x\|^{d}+\|x^{\prime}\|^{d}\right)\left(1+\|x\|^{d_{0}(d+1)}+\|x^{\prime}\|^{d_{0}(d+1)}\right).$
We expand the product as
$(1+\|x\|^{d}+\|x^{\prime}\|^{d})(1+\|x\|^{d_{0}(d+1)}+\|x^{\prime}\|^{d_{0}(d+1)})$
$=1+\|x\|^{d}+\|x^{\prime}\|^{d}+\|x\|^{d_{0}(d+1)}+\|x^{\prime}\|^{d_{0}(d+1)}+\|x\|^{d+d_{0}(d+1)}+\|x^{\prime}\|^{d+d_{0}(d+1)}$
$\quad+\|x\|^{d}\|x^{\prime}\|^{d_{0}(d+1)}+\|x^{\prime}\|^{d}\|x\|^{d_{0}(d+1)}.$
Observe that each of the first seven terms are bounded by $1+\|x\|^{a}+\|x^{\prime}\|^{a}$, where $a$ is given by
$a=d+d_{0}(d+1).$
For the remaining two terms, we apply the weighted AM-GM inequality to obtain
$\|x\|^{d}\|x^{\prime}\|^{d_{0}(d+1)}=(\|x\|^{a})^{\frac{d}{a}}(\|x^{\prime}\|^{a})^{\frac{d_{0}(d+1)}{a}}\leq\frac{d}{a}\|x\|^{a}+\frac{d_{0}(d+1)}{a}\|x^{\prime}\|^{a}\leq\|x\|^{a}+\|x^{\prime}\|^{a}.$
The same bound applies to $\|x^{\prime}\|^{d}\|x\|^{d_{0}(d+1)}$. Therefore the entire product satisfies
$(1+\|x\|^{d}+\|x^{\prime}\|^{d})(1+\|x\|^{d_{0}(d+1)}+\|x^{\prime}\|^{d_{0}(d+1)})\lesssim 1+\|x\|^{a}+\|x^{\prime}\|^{a},$
which implies that $h$ is pseudo-Lipschitz of order $a$. ∎
Lemma 7.
Define $f:\mathbb{R}^{2}\to\mathbb{R}$ by $f(x,y)=xy$. Then $f$ is pseudo-Lipschitz of order $d$ for every $d\in[2,\infty)$.
Proof.
For $(x,y),(x^{\prime},y^{\prime})\in\mathbb{R}^{2}$, we have
$|f(x,y)-f(x^{\prime},y^{\prime})|=|xy-x^{\prime}y^{\prime}|=|x(y-y^{\prime})+(x-x^{\prime})y^{\prime}|$
$\leq|x||y-y^{\prime}|+|x-x^{\prime}||y^{\prime}|\leq\|(x,y)-(x^{\prime},y^{\prime})\|(|x|+|y^{\prime}|).$
Observe that for any $d\geq 2$, we have
$|x|\leq\|(x,y)\|\leq 1+\|(x,y)\|^{d-1},\quad|y^{\prime}|\leq\|(x^{\prime},y^{\prime})\|\leq 1+\|(x^{\prime},y^{\prime})\|^{d-1},$
and consequently, we have
$|x|+|y^{\prime}|\lesssim 1+\|(x,y)\|^{d-1}+\|(x^{\prime},y^{\prime})\|^{d-1}.$
This gives us
$|f(x,y)-f(x^{\prime},y^{\prime})|\lesssim\|(x,y)-(x^{\prime},y^{\prime})\|(1+\|(x,y)\|^{d-1}+\|(x^{\prime},y^{\prime})\|^{d-1}),$
which shows that $f$ is pseudo-Lipschitz of order $d$. ∎
Appendix C Remaining proofs
C.1. Proof of Theorem 3
In this section we provide detailed proofs for the results sketched in Section 4, thereby completing the proof of Theorem 3.
Throughout, $\mathbb{E}[\cdot\mid X]$ denotes the conditional expectation with respect to the $\sigma$-algebra $\sigma(X)$. Since conditional expectations are only defined up to almost-sure equality, we omit “$\mathrm{a.s.}$” when writing “=$\stackrel{{\scriptstyle\mathrm{a.s.}}}{{=}}$” in this context.
C.1.1. Weak Convergence of the Dot Products
In this section we prove the following proposition.
Proposition 8.
Under the assumptions of Theorem 3, the vector $(p_{1},\dots,p_{r})$ converges in distribution to $(\mathring{p}_{1},\dots,\mathring{p}_{r})$, which is the Gaussian vector defined in Definition 1.
The proof of Proposition 8 relies on the next lemma, which is an application of Theorem 2 in BCRT (58).
Lemma 9.
For each $n\in\mathbb{N}$, let $\{X_{\alpha}\}_{\alpha\in[n]}$ be an exchangeable sequence of random variables satisfying
$\mathbb{E}[X_{\alpha}]=0,\quad\mathbb{E}[X_{\alpha}^{2}]=\sigma_{n}^{2},\quad\sigma_{n}^{2}\to\sigma_{*}^{2}\geq 0\quad(n\to\infty).$
Set $S=\sum_{\alpha=1}^{n}X_{\alpha}/\sqrt{n}$. Assume the following conditions:
  1. (a)
    $\mathbb{E}(X_{1}X_{2})=o(1/n)$.
  2. (b)
    $\lim_{n\to\infty}\mathbb{E}(X_{1}^{2}X_{2}^{2})=\sigma_{*}^{4}$.
  3. (c)
    $\mathbb{E}(|X_{1}|^{3})=o(\sqrt{n})$.
Then, $S$ converges in distribution to $Z$, where the random variable $Z$ satisfies
$Z\stackrel{{\scriptstyle\mathrm{a.s.}}}{{=}}0\quad\text{if}\quad\sigma_{*}^{2}=0,\qquad Z\sim N(0,\sigma_{*}^{2})\quad\text{ otherwise}.$
We introduce the following notation. For any two matrices $W^{i,j}$ and $W^{i^{\prime},j^{\prime}}$ appearing in the program, define $d_{(i,j)}^{(i^{\prime},j^{\prime})}$ by
$d_{(i,j)}^{(i^{\prime},j^{\prime})}=\begin{cases}1&(\text{if }W^{i,j}\text{ and }W^{i^{\prime},j^{\prime}}\text{ are the same matrices}),\\ 0&(\text{otherwise}).\end{cases}$
It is important to note that $d_{(i,j)}^{(i^{\prime},j^{\prime})}$ is always a deterministic value that is independ of $n$, and is fixed by the program architecture. According to the sampling rule explained in Section 2.1, the matrices $W^{i,j}$ and $W^{i^{\prime}j^{\prime}}$ are sampled independently whenever $d_{(i,j)}^{(i^{\prime},j^{\prime})}$ is zero. In particular, Assumption 1 gives
$d_{(i,1)}^{(i,2)}=0\quad(i\in[r]).$ (8)
Let $t_{1},\dots,t_{r}$ be arbitrary constants. We define
$S=\sum_{i=1}^{r}t_{i}p_{i}=\sum_{i=1}^{r}t_{i}\left(\frac{1}{\sqrt{n}}\sum_{\alpha=1}^{n}\sum_{\gamma_{1}=1}^{n}\sum_{\gamma_{2}=1}^{n}W_{\alpha\gamma_{1}}^{i,1}W_{\alpha\gamma_{2}}^{i,2}x_{\gamma_{1}}^{i,1}x_{\gamma_{2}}^{i,2}\right)=\frac{1}{\sqrt{n}}\sum_{\alpha=1}^{n}X_{\alpha},$
where $X_{\alpha}$ is given by
$X_{\alpha}=\sum_{i=1}^{r}\sum_{\gamma_{1},\gamma_{2}=1}^{n}t_{i}W_{\alpha\gamma_{1}}^{i,1}W_{\alpha\gamma_{2}}^{i,2}x_{\gamma_{1}}^{i,1}x_{\gamma_{2}}^{i,2}.$
Lemmas 1015 show that the sequence $\{X_{\alpha}\}_{\alpha\in[n]}$ satisfies the conditions of Lemma 9. The proof of Proposition 8 is completed by applying Lemma 9 to $S=\sum_{\alpha=1}^{n}X_{\alpha}/\sqrt{n}$ and then invoking the Cramér–Wold device.
Lemma 10.
The sequence $\{X_{\alpha}\}_{\alpha\in[n]}$ is exchangeable.
Proof.
By the sampling rule, an element $W_{\alpha\beta}^{i,j}$ of the random matrix $W^{i,j}$ independently and identically follows $\mathcal{N}(0,\sigma_{W^{i,j}}^{2}/n)$ for $\alpha,\beta\in[n]$. Hence, conditional on $\{x^{i,j}:i\in[r],\ j\in[2]\}$, the random variables $X_{1},\dots,X_{n}$ are i.i.d. Hence, by de Finetti’s theorem, $\{X_{\alpha}\}_{\alpha\in[n]}$ is exchangeable. ∎
Lemma 11.
$\mathbb{E}(X_{\alpha})=0$ holds for every $\alpha\in[n]$.
Proof.
We compute $\mathbb{E}(X_{\alpha})=\sum_{i=1}^{r}\sum_{\gamma_{1},\gamma_{2}=1}^{n}t_{i}\mathbb{E}\left(W_{\alpha\gamma_{1}}^{i,1}\right)\mathbb{E}\left(W_{\alpha\gamma_{2}}^{i,2}\right)\mathbb{E}\left(x_{\gamma_{1}}^{i,1}x_{\gamma_{2}}^{i,2}\right)=0$. ∎
Lemma 12.
We have $\lim_{n\to\infty}\mathbb{E}(X_{\alpha}^{2})=\sigma_{*}^{2}$ with
$\sigma_{*}^{2}$ $=\sum_{i_{1},i_{2}=1}^{r}t_{i_{1}}t_{i_{2}}\mathbb{E}\left[Z^{g^{i_{1},1}}Z^{g^{i_{1},2}}Z^{g^{i_{2},1}}Z^{g^{i_{2},2}}\right]$
$=\sum_{i_{1},i_{2}=1}^{r}\sum_{(j,j^{\prime})\in J}t_{i_{1}}t_{i_{2}}\sigma_{W^{i_{1},1}}^{2}\sigma_{W^{i_{1},2}}^{2}\mathbb{E}\left[Z^{x^{i_{1},1}}Z^{x^{i_{2},j}}\right]\mathbb{E}\left[Z^{x^{i_{1},2}}Z^{x^{i_{2},j^{\prime}}}\right]d_{(i_{1},1)}^{(i_{2},j)}d_{(i_{1},2)}^{(i_{2},j^{\prime})},$
where we defined $J=\{(1,2),(2,1)\}$.
Proof.
For any $\alpha\in[n]$, we have
$\mathbb{E}(X_{\alpha}^{2})$
$=\mathbb{E}\left[\left(\sum_{i_{1}=1}^{r}\sum_{\gamma_{1},\gamma_{2}=1}^{n}t_{i_{1}}W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{1},2}x_{\gamma_{1}}^{i,1}x_{\gamma_{2}}^{i,2}\right)\left(\sum_{i_{2}=1}^{r}\sum_{\gamma_{3},\gamma_{4}=1}^{n}t_{i_{2}}W_{\alpha\gamma_{3}}^{i_{2},1}W_{\alpha\gamma_{4}}^{i_{2},2}x_{\gamma_{3}}^{i_{2},1}x_{\gamma_{4}}^{i_{2},2}\right)\right]$
$=\sum_{i_{1},i_{2}=1}^{r}\sum_{\gamma_{1},\dots,\gamma_{4}=1}^{n}t_{i_{1}}t_{i_{2}}\mathbb{E}\left(W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{1},2}W_{\alpha\gamma_{3}}^{i_{2},1}W_{\alpha\gamma_{4}}^{i_{2},2}\right)\mathbb{E}\left(x_{\gamma_{1}}^{i_{1},1}x_{\gamma_{2}}^{i_{1},2}x_{\gamma_{3}}^{i_{2},1}x_{\gamma_{4}}^{i_{2},2}\right)$
$=\sum_{i_{1},i_{2}=1}^{r}\sum_{\gamma_{1},\gamma_{2}=1}^{n}\sum_{(j,j^{\prime})\in J}t_{i_{1}}t_{i_{2}}\mathbb{E}\left[(W_{\alpha\gamma_{1}}^{i_{1},1})^{2}\right]\mathbb{E}\left[(W_{\alpha\gamma_{2}}^{i_{1},2})^{2}\right]\mathbb{E}\left(x_{\gamma_{1}}^{i_{1},1}x_{\gamma_{1}}^{i_{2},j}x_{\gamma_{2}}^{i_{1},2}x_{\gamma_{2}}^{i_{2},j^{\prime}}\right)d_{(i_{1},1)}^{(i_{2},j)}d_{(i_{1},2)}^{(i_{2},j^{\prime})}$
$=\sum_{i_{1},i_{2}=1}^{r}\sum_{(j,j^{\prime})\in J}t_{i_{1}}t_{i_{2}}\sigma_{W^{i_{1},1}}^{2}\sigma_{W^{i_{1},2}}^{2}\mathbb{E}\left[\left(\frac{1}{n}\sum_{\gamma_{1}=1}^{n}x_{\gamma_{1}}^{i_{1},1}x_{\gamma_{1}}^{i_{2},j}\right)\left(\frac{1}{n}\sum_{\gamma_{2}=1}^{n}x_{\gamma_{2}}^{i_{1},2}x_{\gamma_{2}}^{i_{2},j^{\prime}}\right)\right]d_{(i_{1},1)}^{(i_{2},j)}d_{(i_{1},2)}^{(i_{2},j^{\prime})}$
$\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}\sum_{i_{1},i_{2}=1}^{r}\sum_{(j,j^{\prime})\in J}t_{i_{1}}t_{i_{2}}\sigma_{W^{i_{1},1}}^{2}\sigma_{W^{i_{1},2}}^{2}\mathbb{E}\left[Z^{x^{i_{1},1}}Z^{x^{i_{2},j}}\right]\mathbb{E}\left[Z^{x^{i_{1},2}}Z^{x^{i_{2},j^{\prime}}}\right]d_{(i_{1},1)}^{(i_{2},j)}d_{(i_{1},2)}^{(i_{2},j^{\prime})},$
where the convergence follows from Lemma 16. Finally, Eq. (8) and Definition 1 imply that
$\sigma_{W^{i_{1},1}}^{2}\sigma_{W^{i_{1},2}}^{2}\mathbb{E}\left[Z^{x^{i_{1},1}}Z^{x^{i_{2},j}}\right]\mathbb{E}\left[Z^{x^{i_{1},2}}Z^{x^{i_{2},j^{\prime}}}\right]d_{(i_{1},1)}^{(i_{2},j)}d_{(i_{1},2)}^{(i_{2},j^{\prime})}=\mathbb{E}\left[Z^{g^{i_{1},1}}Z^{g^{i_{1},2}}Z^{g^{i_{2},1}}Z^{g^{i_{2},2}}\right]$
holds for any $i_{1},i_{2}\in[r]$ and $(j,j^{\prime})\in J$. ∎
Lemma 13.
$\mathbb{E}(X_{\alpha}X_{\beta})=0$ holds for every $\alpha,\beta\in[n],\alpha\neq\beta$.
Proof.
For $\alpha\neq\beta$, we have
$\mathbb{E}(X_{\alpha}X_{\beta})$ $=\mathbb{E}\left[\left(\sum_{i_{1}=1}^{r}\sum_{\gamma_{1},\gamma_{2}=1}^{n}t_{i_{1}}W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{1},2}x_{\gamma_{1}}^{i_{1},1}x_{\gamma_{2}}^{i_{1},2}\right)\left(\sum_{i_{2}=1}^{r}\sum_{\gamma_{3},\gamma_{4}=1}^{n}t_{i_{2}}W_{\beta\gamma_{3}}^{i_{2},1}W_{\beta\gamma_{4}}^{i_{2},2}x_{\gamma_{3}}^{i_{2},1}x_{\gamma_{4}}^{i_{2},2}\right)\right]$
$=\sum_{i_{1},i_{2}=1}^{r}\sum_{\gamma_{1},\dots,\gamma_{4}=1}^{n}t_{i_{1}}t_{i_{2}}\mathbb{E}\left(W_{\alpha\gamma_{1}}^{i_{1},1}\right)\mathbb{E}\left(W_{\alpha\gamma_{2}}^{i_{1},2}\right)\mathbb{E}\left(W_{\beta\gamma_{3}}^{i_{2},1}\right)\mathbb{E}\left(W_{\beta\gamma_{4}}^{i_{2},2}\right)\mathbb{E}\left(x_{\gamma_{1}}^{i_{1},1}x_{\gamma_{2}}^{i_{1},2}x_{\gamma_{3}}^{i_{2},1}x_{\gamma_{4}}^{i_{2},2}\right)$
$=0$
as desired. ∎
Lemma 14.
$\lim_{n\to\infty}\mathbb{E}(X_{\alpha}^{2}X_{\beta}^{2})=\sigma_{*}^{4}$ holds for every $\alpha,\beta\in[n],\alpha\neq\beta$.
Proof.
By a calculation similar to Lemma 12, we have
$\mathbb{E}(X_{\alpha}^{2}X_{\beta}^{2})$
$=\sum_{i_{1},\dots,i_{4}=1}^{r}\sum_{(j_{1},j_{1}^{\prime}),(j_{2},j_{2}^{\prime})\in J^{2}}t_{i_{1}}t_{i_{2}}t_{i_{3}}t_{i_{4}}\sigma_{W^{i_{1},1}}^{2}\sigma_{W^{i_{2},2}}^{2}\sigma_{W^{i_{3},1}}^{2}\sigma_{W^{i_{4},2}}^{2}d_{(i_{1},1)}^{(i_{2},j_{1})}d_{(i_{1},2)}^{(i_{2},j_{1}^{\prime})}d_{(i_{3},1)}^{(i_{4},j_{2})}d_{(i_{3},2)}^{(i_{4},j_{2}^{\prime})}$
×$\quad\times\mathbb{E}\left[\left(\frac{1}{n}\sum_{\gamma_{1}=1}^{n}x_{\gamma_{1}}^{i_{1},1}x_{\gamma_{1}}^{i_{2},j_{1}}\right)\left(\frac{1}{n}\sum_{\gamma_{2}=1}^{n}x_{\gamma_{2}}^{i_{1},2}x_{\gamma_{2}}^{i_{2},j_{1}^{\prime}}\right)\left(\frac{1}{n}\sum_{\gamma_{3}=1}^{n}x_{\gamma_{3}}^{i_{3},1}x_{\gamma_{3}}^{i_{4},j_{2}}\right)\left(\frac{1}{n}\sum_{\gamma_{4}=1}^{n}x_{\gamma_{4}}^{i_{3},2}x_{\gamma_{4}}^{i_{4},j_{2}^{\prime}}\right)\right]$
$\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}\sum_{i_{1},\dots,i_{4}=1}^{r}\sum_{(j_{1},j_{1}^{\prime}),(j_{2},j_{2}^{\prime})\in J^{2}}t_{i_{1}}t_{i_{2}}t_{i_{3}}t_{i_{4}}\sigma_{W^{i_{1},1}}^{2}\sigma_{W^{i_{2},2}}^{2}\sigma_{W^{i_{3},1}}^{2}\sigma_{W^{i_{4},2}}^{2}d_{(i_{1},1)}^{(i_{2},j_{1})}d_{(i_{1},2)}^{(i_{2},j_{1}^{\prime})}d_{(i_{3},1)}^{(i_{4},j_{2})}d_{(i_{3},2)}^{(i_{4},j_{2}^{\prime})}$
×$\quad\times\mathbb{E}\left[Z^{x^{i_{1},1}}Z^{x^{i_{2},j_{1}}}\right]\mathbb{E}\left[Z^{x^{i_{1},2}}Z^{x^{i_{2},j_{1}^{\prime}}}\right]\mathbb{E}\left[Z^{x^{i_{3},1}}Z^{x^{i_{4},j_{2}}}\right]\mathbb{E}\left[Z^{x^{i_{3},2}}Z^{x^{i_{4},j_{2}^{\prime}}}\right],$
where the convergence follows from Lemma 16. Observe that this limit is equivalent to $\sigma_{*}^{4}$. ∎
Lemma 15.
$\mathbb{E}(|X_{\alpha}|^{3})=o(\sqrt{n})$ holds as $n\to\infty$ for every $\alpha\in[n]$.
Proof.
By the Lyapunov inequality, we have
$\frac{1}{\sqrt{n}}\mathbb{E}(|X_{\alpha}|^{3})\leq\frac{1}{\sqrt{n}}\left(\mathbb{E}(X_{\alpha}^{4})\right)^{\frac{3}{4}}\leq\frac{1}{\sqrt{n}}\left(\sup_{n}\mathbb{E}(X_{\alpha}^{4})\right)^{\frac{3}{4}}.$
Thus, it suffices to show that $\sup_{n}\mathbb{E}(X_{\alpha}^{4})<\infty$ holds. We can express $\mathbb{E}(X_{\alpha}^{4})$ as
$\mathbb{E}(X_{\alpha}^{4})$ $=\sum_{i_{1},\dots,i_{4}=1}^{r}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}t_{i_{1}}t_{i_{2}}t_{i_{3}}t_{i_{4}}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{2},1}W_{\alpha\gamma_{3}}^{i_{3},1}W_{\alpha\gamma_{4}}^{i_{4},1}W_{\alpha\gamma_{5}}^{i_{1},2}W_{\alpha\gamma_{6}}^{i_{2},2}W_{\alpha\gamma_{7}}^{i_{3},2}W_{\alpha\gamma_{8}}^{i_{4},2}\right]$
×$\quad\times\mathbb{E}\left[x_{\gamma_{1}}^{i_{1},1}x_{\gamma_{2}}^{i_{2},1}x_{\gamma_{3}}^{i_{3},1}x_{\gamma_{4}}^{i_{4},1}x_{\gamma_{5}}^{i_{1},2}x_{\gamma_{6}}^{i_{2},2}x_{\gamma_{7}}^{i_{3},2}x_{\gamma_{8}}^{i_{4},2}\right].$
Applying the Cauchy–Schwarz inequality yields
$\mathbb{E}\left[x_{\gamma_{1}}^{i_{1},1}x_{\gamma_{2}}^{i_{2},1}x_{\gamma_{3}}^{i_{3},1}x_{\gamma_{4}}^{i_{4},1}x_{\gamma_{5}}^{i_{1},2}x_{\gamma_{6}}^{i_{2},2}x_{\gamma_{7}}^{i_{3},2}x_{\gamma_{8}}^{i_{4},2}\right]$
$\leq\left(\prod_{k=1}^{4}\mathbb{E}\left[(x_{\gamma_{j}}^{i_{j},1})^{8}\right]\right)^{\frac{1}{8}}\left(\prod_{k=1}^{4}\mathbb{E}\left[(x_{\gamma_{4+j}}^{i_{j},2})^{8}\right]\right)^{\frac{1}{8}}=\left(\prod_{k=1}^{4}\mathbb{E}\left[(x_{1}^{i_{j},1})^{8}\right]\right)^{\frac{1}{8}}\left(\prod_{k=1}^{4}\mathbb{E}\left[(x_{1}^{i_{j},2})^{8}\right]\right)^{\frac{1}{8}}$
$\leq\sup_{i\in[r],\ j\in[2]}\mathbb{E}\left[(x_{1}^{i,j})^{8}\right].$
By the boundedness of $x^{i,j}\ (i\in[r],\ j\in[2])$, the last term is bounded uniformly in $n$, and consequently, it holds that
$\sup_{n}\mathbb{E}\left[x_{\gamma_{1}}^{i_{1},1}x_{\gamma_{2}}^{i_{2},1}x_{\gamma_{3}}^{i_{3},1}x_{\gamma_{4}}^{i_{4},1}x_{\gamma_{5}}^{i_{1},2}x_{\gamma_{6}}^{i_{2},2}x_{\gamma_{7}}^{i_{3},2}x_{\gamma_{8}}^{i_{4},2}\right]<\infty.$
Hence, we compute
$\mathbb{E}(X_{\alpha}^{4})$ $\lesssim\sum_{i_{1},\dots,i_{4}=1}^{r}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{2},1}W_{\alpha\gamma_{3}}^{i_{3},1}W_{\alpha\gamma_{4}}^{i_{4},1}W_{\alpha\gamma_{5}}^{i_{1},2}W_{\alpha\gamma_{6}}^{i_{2},2}W_{\alpha\gamma_{7}}^{i_{3},2}W_{\alpha\gamma_{8}}^{i_{4},2}\right]$
$=\sum_{i=1}^{r}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i,1}W_{\alpha\gamma_{2}}^{i,1}W_{\alpha\gamma_{3}}^{i,1}W_{\alpha\gamma_{4}}^{i,1}W_{\alpha\gamma_{5}}^{i,2}W_{\alpha\gamma_{6}}^{i,2}W_{\alpha\gamma_{7}}^{i,2}W_{\alpha\gamma_{8}}^{i,2}\right]$
$\quad+4\sum_{i_{1}=1}^{r}\sum_{i_{2}\neq i_{1}}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{1},1}W_{\alpha\gamma_{3}}^{i_{1},1}W_{\alpha\gamma_{4}}^{i_{2},1}W_{\alpha\gamma_{5}}^{i_{1},2}W_{\alpha\gamma_{6}}^{i_{1},2}W_{\alpha\gamma_{7}}^{i_{1},2}W_{\alpha\gamma_{8}}^{i_{2},2}\right]$
$\quad+3\sum_{i_{1}=1}^{r}\sum_{i_{2}\neq i_{1}}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{1},1}W_{\alpha\gamma_{3}}^{i_{2},1}W_{\alpha\gamma_{4}}^{i_{2},1}W_{\alpha\gamma_{5}}^{i_{1},2}W_{\alpha\gamma_{6}}^{i_{1},2}W_{\alpha\gamma_{7}}^{i_{2},2}W_{\alpha\gamma_{8}}^{i_{2},2}\right]$
$\quad+6\sum_{i_{1}=1}^{r}\sum_{i_{2}\neq i_{1}}\sum_{i_{3}\notin\{i_{1},i_{2}\}}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{1},1}W_{\alpha\gamma_{3}}^{i_{2},1}W_{\alpha\gamma_{4}}^{i_{3},1}W_{\alpha\gamma_{5}}^{i_{1},2}W_{\alpha\gamma_{6}}^{i_{1},2}W_{\alpha\gamma_{7}}^{i_{2},2}W_{\alpha\gamma_{8}}^{i_{3},2}\right]$
$\quad+\sum_{\begin{subarray}{c}i_{1},i_{2},i_{3},i_{4}=1\\ \text{all distinct}\end{subarray}}^{r}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{2},1}W_{\alpha\gamma_{3}}^{i_{3},1}W_{\alpha\gamma_{4}}^{i_{4},1}W_{\alpha\gamma_{5}}^{i_{1},2}W_{\alpha\gamma_{6}}^{i_{2},2}W_{\alpha\gamma_{7}}^{i_{3},2}W_{\alpha\gamma_{8}}^{i_{4},2}\right]$
$=:A_{1}+4A_{2}+3A_{3}+6A_{4}+A_{5}.$
Define $\sigma$ by $\sigma=\max\{\sigma_{W^{i,j}}:i\in[r],\ j\in[2]\}$. Then, we compute $A_{1}$ as
$A_{1}$ $=\sum_{i=1}^{r}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i,1}W_{\alpha\gamma_{2}}^{i,1}W_{\alpha\gamma_{3}}^{i,1}W_{\alpha\gamma_{4}}^{i,1}\right]\mathbb{E}\left[W_{\alpha\gamma_{5}}^{i,2}W_{\alpha\gamma_{6}}^{i,2}W_{\alpha\gamma_{7}}^{i,2}W_{\alpha\gamma_{8}}^{i,2}\right]$
$=\sum_{i=1}^{r}\sum_{\gamma_{1},\gamma_{2}=1}^{n}\mathbb{E}\left[(W_{\alpha\gamma_{1}}^{i,1})^{4}\right]\mathbb{E}\left[(W_{\alpha\gamma_{2}}^{i,2})^{2}\right]$
$\quad+3\sum_{i=1}^{r}\sum_{(j,j^{\prime})\in J}\sum_{\gamma_{1},\gamma_{2}=1}^{n}\sum_{\gamma_{3}\neq\gamma_{2}}\mathbb{E}\left[(W_{\alpha\gamma_{1}}^{i,j})^{4}\right]\mathbb{E}\left[(W_{\alpha\gamma_{2}}^{i,j^{\prime}})^{2}\right]\mathbb{E}\left[(W_{\alpha\gamma_{3}}^{i,j^{\prime}})^{2}\right]$
$\quad+9\sum_{i=1}^{r}\sum_{\gamma_{1},\gamma_{2}=1}^{n}\sum_{\gamma_{3}\neq\gamma_{1}}\sum_{\gamma_{4}\neq\gamma_{2}}\mathbb{E}\left[(W_{\alpha\gamma_{1}}^{i,1})^{2}\right]\mathbb{E}\left[(W_{\alpha\gamma_{2}}^{i,1})^{2}\right]\mathbb{E}\left[(W_{\alpha\gamma_{3}}^{i,2})^{2}\right]\mathbb{E}\left[(W_{\alpha\gamma_{4}}^{i,2})^{2}\right]$
$=\sum_{i=1}^{r}\left(\frac{9\sigma_{W^{i,1}}^{4}\sigma_{W^{i,2}}^{4}}{n^{2}}+3\sum_{(j,j^{\prime})\in J}\frac{3(n-1)\sigma_{W^{i,j}}^{4}\sigma_{W^{i,j^{\prime}}}^{4}}{n^{2}}+9\frac{(n-1)^{2}\sigma_{W^{i,1}}^{4}\sigma_{W^{i,2}}^{4}}{n^{2}}\right)$
$=9\sum_{i=1}^{r}\sigma_{W^{i,1}}^{4}\sigma_{W^{i,2}}^{4}\lesssim\sigma^{8}.$
Likewise, we have
$A_{2}$ $=\sum_{i_{1}=1}^{r}\sum_{i_{2}\neq i_{1}}\sum_{(j,j^{\prime})\in J}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{1},1}W_{\alpha\gamma_{3}}^{i_{1},1}W_{\alpha\gamma_{4}}^{i_{2},j}\right]\mathbb{E}\left[W_{\alpha\gamma_{5}}^{i_{1},2}W_{\alpha\gamma_{6}}^{i_{1},2}W_{\alpha\gamma_{7}}^{i_{1},2}W_{\alpha\gamma_{8}}^{i_{2},j^{\prime}}\right]d_{(i_{1},1)}^{(i_{2},j)}d_{(i_{1},2)}^{(i_{2},j^{\prime})}$
$=9\sum_{i_{1}=1}^{r}\sum_{i_{2}\neq i_{1}}\sum_{(j,j^{\prime})\in J}\sigma_{W^{i_{1},1}}^{4}\sigma_{W^{i_{1},2}}^{4}d_{(i_{1},1)}^{(i_{2},j)}d_{(i_{1},2)}^{(i_{2},j^{\prime})}\lesssim\sigma^{8}$
and
$A_{3}$ $=\sum_{i_{1}=1}^{r}\sum_{i_{2}\neq i_{1}}\sum_{(j,j^{\prime})\in J}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{1},1}W_{\alpha\gamma_{3}}^{i_{2},j}W_{\alpha\gamma_{4}}^{i_{2},j}\right]\mathbb{E}\left[W_{\alpha\gamma_{5}}^{i_{1},2}W_{\alpha\gamma_{6}}^{i_{1},2}W_{\alpha\gamma_{7}}^{i_{2},j^{\prime}}W_{\alpha\gamma_{8}}^{i_{2},j^{\prime}}\right]d_{(i_{1},1)}^{(i_{2},j)}d_{(i_{1},2)}^{(i_{2},j^{\prime})}$
$=9\sum_{i_{1}=1}^{r}\sum_{i_{2}\neq i_{1}}\sum_{(j,j^{\prime})\in J}\sigma_{W^{i_{1},1}}^{4}\sigma_{W^{i_{1},2}}^{4}d_{(i_{1},1)}^{(i_{2},j)}d_{(i_{1},2)}^{(i_{2},j^{\prime})}\lesssim\sigma^{8}.$
Applying a similar argument, we can also show that $A_{4}\lesssim\sigma^{8}$ and $A_{5}\lesssim\sigma^{8}$ holds. Therefore, we conclude that $\mathbb{E}(X_{\alpha}^{4})\lesssim\sigma^{8}$ holds, which completes the proof. ∎
Lemma 16.
Take $k_{1},\dots,k_{8}\in\{(i,j):i\in[r],\ j\in[2]\}$ arbitrarily. Then, the following statements hold as $n\to\infty$.
  1. (i)
    For $(x^{k_{1}},\dots,x^{k_{8}})$, we have
    $\mathbb{E}\left[\left(\frac{1}{n}\sum_{\gamma_{1}=1}^{n}x_{\gamma_{1}}^{k_{1}}x_{\gamma_{1}}^{k_{2}}\right)\left(\frac{1}{n}\sum_{\gamma_{2}=1}^{n}x_{\gamma_{2}}^{k_{3}}x_{\gamma_{2}}^{k_{4}}\right)\left(\frac{1}{n}\sum_{\gamma_{3}=1}^{n}x_{\gamma_{3}}^{k_{5}}x_{\gamma_{3}}^{k_{6}}\right)\left(\frac{1}{n}\sum_{\gamma_{4}=1}^{n}x_{\gamma_{4}}^{k_{7}}x_{\gamma_{4}}^{k_{8}}\right)\right]$
    $\to\mathbb{E}\left[Z^{x^{k_{1}}}Z^{x^{k_{2}}}\right]\mathbb{E}\left[Z^{x^{k_{3}}}Z^{x^{k_{4}}}\right]\mathbb{E}\left[Z^{x^{k_{5}}}Z^{x^{k_{6}}}\right]\mathbb{E}\left[Z^{x^{k_{7}}}Z^{x^{k_{8}}}\right].$
  2. (ii)
    For $(x^{k_{1}},\dots,x^{k_{4}})$, we have
    $\mathbb{E}\left[\left(\frac{1}{n}\sum_{\gamma_{1}=1}^{n}x_{\gamma_{1}}^{k_{1}}x_{\gamma_{1}}^{k_{2}}\right)\left(\frac{1}{n}\sum_{\gamma_{2}=1}^{n}x_{\gamma_{2}}^{k_{3}}x_{\gamma_{2}}^{k_{4}}\right)\right]\to\mathbb{E}\left[Z^{x^{k_{1}}}Z^{x^{k_{2}}}\right]\mathbb{E}\left[Z^{x^{k_{3}}}Z^{x^{k_{4}}}\right].$
Proof.
Define the residual term $R_{\ell}$ by
$R_{\ell}=\frac{1}{n}\sum_{\gamma_{1}=1}^{n}x_{\gamma_{\ell}}^{k_{2_{\ell}-1}}x_{\gamma_{\ell}}^{k_{2_{\ell}}}-\mathbb{E}\left[Z^{x^{k_{2_{\ell}-1}}}Z^{x^{k_{2_{\ell}}}}\right]$
for each $i\in[4]$. Define constants $C$ and $\tilde{R}$ by
$C=\max_{\ell\in[4]}\left|\mathbb{E}\left[Z^{x^{k_{2_{\ell}-1}}}Z^{x^{k_{2_{\ell}}}}\right]\right|,\qquad\tilde{R}=\left(\max_{\ell\in[4]}\mathbb{E}(R_{\ell}^{4})\right)^{\frac{1}{4}}.$
Note that $C$ is bounded by the boundedness of $x^{i,j}$ for all $i\in[r]$ and $j\in[2]$ (see Definition 1). Then, we can write
$\left|\mathbb{E}\left[\prod_{\ell=1}^{4}\left(\frac{1}{n}\sum_{\gamma_{\ell}=1}^{n}x_{\gamma_{\ell}}^{k_{2_{\ell}-1}}x_{\gamma_{\ell}}^{k_{2_{\ell}}}\right)\right]-\prod_{\ell=1}^{4}\mathbb{E}\left[Z^{x^{k_{2_{\ell}-1}}}Z^{x^{k_{2_{\ell}}}}\right]\right|$
$=\left|\mathbb{E}\left[\prod_{\ell=1}^{4}\left(R_{\ell}+\mathbb{E}\left[Z^{x^{k_{2_{\ell}-1}}}Z^{x^{k_{2_{\ell}}}}\right]\right)\right]-\prod_{\ell=1}^{4}\mathbb{E}\left[Z^{x^{k_{2_{\ell}-1}}}Z^{x^{k_{2_{\ell}}}}\right]\right|$
$\leq 4C^{3}\tilde{R}+6C^{2}\tilde{R}^{2}+4C\tilde{R}^{3}+\tilde{R}^{4},$
where the last inequality follows from the Cauchy–Schwarz inequality and the Lyapunov inequality. Likewise, we have
$\left|\mathbb{E}\left[\prod_{\ell=1}^{2}\left(\frac{1}{n}\sum_{\gamma_{\ell}=1}^{n}x_{\gamma_{\ell}}^{k_{2_{\ell}-1}}x_{\gamma_{\ell}}^{k_{2_{\ell}}}\right)\right]-\prod_{\ell=1}^{2}\mathbb{E}\left[Z^{x^{k_{2_{\ell}-1}}}Z^{x^{k_{2_{\ell}}}}\right]\right|$
$=\left|\mathbb{E}\left[\prod_{\ell=1}^{2}\left(R_{\ell}+\mathbb{E}\left[Z^{x^{k_{2_{\ell}-1}}}Z^{x^{k_{2_{\ell}}}}\right]\right)\right]-\prod_{\ell=1}^{2}\mathbb{E}\left[Z^{x^{k_{2_{\ell}-1}}}Z^{x^{k_{2_{\ell}}}}\right]\right|$
$\leq 2C\tilde{R}+\tilde{R}^{2}.$
Thus, it remains only to prove that $\tilde{R}$ converges to 00 as $n\to\infty$. This can be achieved by showing that $\mathbb{E}(R_{\ell}^{4})$ converges to 00 for all $\ell\in[4]$. By Fact 1 and Lemma 7, for all $\ell\in[4]$, we know $R_{\ell}$ converges almost surely to 00. The continuous mapping theorem then yields
$R_{\ell}^{4}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0.$
To upgrade this to convergence in expectation, Facts 2 and 3 imply it suffices to show the existance of a constant $\delta>1$ that satisfies
$\sup_{n}\mathbb{E}(R_{\ell}^{4+\delta})<\infty.$
But since each $x^{i,j}$ and its infinite-width limit $Z^{x^{i,j}}$ are bounded, such a $\delta$ exists. Therefore, we conclude that $\mathbb{E}(R_{\ell}^{4})$ converges to 00 for all $\ell\in[4]$, and consequently, $\tilde{R}$ does as well. ∎
C.1.2. $S_{1}$ Converges to 0
We study the convergence of the term $S_{1}$ defined in Section 4. Specifically, we show
$S_{1}=\left|\mathbb{E}f\left(\frac{1}{n}\sum_{\alpha=1}^{n}\psi(\bm{g}_{\alpha}^{1:M},\bm{p}_{1:r})\right)-\mathbb{E}f\left(\mathbb{E}\left[\psi(Z^{\bm{g}^{1:M}},\bm{p}_{1:r})\mid\bm{p}_{1:r}\right]\right)\right|\to 0.$
Fix a small $\epsilon>0$. Since $\mathring{\bm{p}}_{1:r}$ is a Gaussian vector (see Definition 1), it is tight. Hence there exists a compact set $K\subset\mathbb{R}^{r}$ that satisfies
$\mathrm{P}(\mathring{\bm{p}}_{1:r}\in K)>1-\epsilon.$
Set
$y=\frac{1}{n}\sum_{\alpha=1}^{n}\psi(\bm{g}_{\alpha}^{1:M},\bm{p}_{1:r}),\quad z=\mathbb{E}\left[\psi(Z^{\bm{g}^{1:M}},\bm{p}_{1:r})\mid\bm{p}_{1:r}\right].$
Then, we have
$S_{1}$ $=\left|\mathbb{E}f(y)-\mathbb{E}f(z)\right|$
$\leq\left|\mathbb{E}[(f(y)-f(z))1\{\bm{p}_{1:r}\in K\}]\right|+\left|\mathbb{E}[(f(y)-f(z))1\{\bm{p}_{1:r}\notin K\}]\right|$
$\leq\left|\mathbb{E}[(f(y)-f(z))1\{\bm{p}_{1:r}\in K\}]\right|+\mathbb{E}[|f(y)-f(z)|1\{\bm{p}_{1:r}\notin K\}].$ (9)
By the boundedness of $f$, the second term of Eq. (9) is bounded as
$\mathbb{E}[|f(y)-f(z)|1\{\bm{p}_{1:r}\notin K\}]\leq C\mathrm{P}(\bm{p}_{1:r}\notin K)$
with some constant $C$. Furthermore, noting that $K$ is a Borel set, applying Lemma 5 and Proposition 8 gives
$\lim_{n\to\infty}\mathrm{P}(\bm{p}_{1:r}\notin K)=\mathrm{P}(\mathring{\bm{p}}_{1:r}\notin K)<\epsilon.$
Therefore, for large enough $n$, we have
$\mathbb{E}[|f(y)-f(z)|1\{\bm{p}_{1:r}\notin K\}]<C\epsilon.$
For the first term of Eq. (9), we have
$\left|\mathbb{E}[(f(y)-f(z))1\{\bm{p}_{1:r}\in K\}]\right|$
$\leq\sup_{\bm{a}_{1:r}\in K}\left|\mathbb{E}\left[f\left(\frac{1}{n}\sum_{\alpha=1}^{n}\psi(\bm{g}_{\alpha}^{1:M},\bm{a}_{1:r})\right)\right]-\mathbb{E}\left[f\left(\mathbb{E}\left[\psi(Z^{\bm{g}^{1:M}},\bm{a}_{1:r})\right]\right)\right]\right|$
$=\sup_{\bm{a}_{1:r}\in K}\Phi_{n}(\bm{a}_{1:r}),$
where $\Phi_{n}:\mathbb{R}^{r}\to\mathbb{R}$ is defined by
$\Phi_{n}(\bm{a}_{1:r})=\left|\mathbb{E}\left[f\left(\frac{1}{n}\sum_{\alpha=1}^{n}\psi(\bm{g}_{\alpha}^{1:M},\bm{a}_{1:r})\right)\right]-\mathbb{E}\left[f\left(\mathbb{E}\left[\psi(Z^{\bm{g}^{1:M}},\bm{a}_{1:r})\right]\right)\right]\right|.$ (10)
Define $\tilde{\psi}_{\bm{a}_{1:r}}:\mathbb{R}^{M}\to\mathbb{R}$ by
$\tilde{\psi}_{\bm{a}_{1:r}}(\bm{u}_{1:M})=\psi(\bm{u}_{1:M},\bm{a}_{1:r}).$ (11)
Observe that $\tilde{\psi}_{\bm{a}_{1:r}}$ is bounded by the boundedness of $\psi$. We later show that $\tilde{\psi}_{\bm{a}_{1:r}}$ is also pseudo-Lipschitz (Lemma 17). Therefore, by Eq. (7) and Lemma 5, we have
$\lim_{n\to\infty}\Phi_{n}(\bm{a}_{1:r})=0\quad(\bm{a}_{1:r}\in\mathbb{R}^{r}).$
Moreover, noting that $f$ and $\psi$ are bounded and continuous (see Fact 4), we can apply the (uncoditional and conditional) bounded convergence theorem to show that $\Phi_{n}$ is continuous at every point. Therefore, we can apply Lemma 18 to show that $\sup_{\bm{a}_{1:r}\in K}\Phi_{n}(\bm{a}_{1:r})$ converges to 00 as $n\to\infty$.
Lemma 17.
Define $\tilde{\psi}_{\bm{a}_{1:r}}:\mathbb{R}^{M}\to\mathbb{R}$ by Eq. (11). Then, it is pseudo-Lipschitz.
Proof.
Suppose $\psi$ is pseudo-Lipschitz of order $d+1$ with $d\geq 1$. Then, we have
$|\tilde{\psi}_{\bm{a}_{1:r}}(\bm{u}_{1:M})-\tilde{\psi}_{\bm{a}_{1:r}}(\bm{u}_{1:M}^{\prime})|=|\psi(\bm{u}_{1:M},\bm{a}_{1:r})-\psi(\bm{u}_{1:M}^{\prime},\bm{a}_{1:r})|$
$\lesssim\|(\bm{u}_{1:M})-(\bm{u}_{1:M}^{\prime})\|(1+\|(\bm{u}_{1:M},\bm{a}_{1:r})\|^{d}+\|(\bm{u}_{1:M}^{\prime},\bm{a}_{1:r})\|^{d}).$
By Lemma 4, we can bound $\|(\bm{u}_{1:M},\bm{a}_{1:r})\|^{d}$ as
$\|(\bm{u}_{1:M},\bm{a}_{1:r})\|^{d}\leq(\|(\bm{u}_{1:M})\|+\|\bm{a}_{1:r}\|)^{d}\leq 2^{d-1}\left(\|(\bm{u}_{1:M})\|^{d}+\|\bm{a}_{1:r}\|^{d}\right).$
Thus, we have
$1+\|(\bm{u}_{1:M},\bm{a}_{1:r})\|^{d}+\|(\bm{u}_{1:M}^{\prime},\bm{a}_{1:r})\|^{d}$ $\leq 1+2^{d-1}\left(\|(\bm{u}_{1:M})\|^{d}+\|(\bm{u}_{1:M}^{\prime})\|^{d}\right)+2^{d}\|\bm{a}_{1:r}\|^{d}$
$\lesssim 1+\|(\bm{u}_{1:M})\|^{d}+\|(\bm{u}_{1:M}^{\prime})\|^{d},$
where the last inequality holds because $\bm{a}_{1:r}$ is fixed. ∎
Lemma 18.
Let $K\subset\mathbb{R}^{r}$ be a compact set. Suppose for each $n\in\mathbb{N}$, $f_{n}:\mathbb{R}\to\mathbb{R}$ is a continuous function that satisfies
$\lim_{n\to\infty}f_{n}(\bm{a}_{1:r})=0$ (12)
for any constant $\bm{a}_{1:r}\in K$. Then, we have
$\lim_{n\to\infty}\sup_{\bm{a}_{1:r}\in K}|f_{n}(\bm{a}_{1:r})|=0.$
Proof.
Fix $\epsilon>0$. Let $B(\bm{a}_{1:r},\epsilon)$ denote the ball of radius $\epsilon$ centered at $\bm{a}_{1:r}\in\mathbb{R}^{r}$. By the continuity of $f_{n}$, for each $\bm{a}_{1:r}\in K$, there exists $\delta_{\bm{a}_{1:r}}>0$ such that
$|f_{n}(\bm{a}_{1:r})-f_{n}(\bm{b}_{1:r})|<\epsilon/2\quad(\bm{b}_{1:r}\in B(\bm{a}_{1:r},\delta_{\bm{a}_{1:r}}))$
holds. Note that $K$ is covered by the union of balls as
$K\subset\bigcup_{\bm{a}_{1:r}\in K}B(\bm{a}_{1:r},\delta_{\bm{a}_{1:r}}).$
Since $K$ is compact, we can cover $K$ with finitely many balls, say
$K\subset\bigcup_{i=1}^{I}B_{i}=\bigcup_{i=1}^{I}B(\bm{a}_{1:r}^{i},\delta_{\bm{a}_{1:r}^{i}}),$
where $B_{i}$ is given by $B_{i}=B(\bm{a}_{1:r}^{i},\delta_{a^{i}})$. By Eq. (12), for each $i\in[I]$, there exists $N_{i}\in\mathbb{N}$ such that
$|f_{n}(\bm{a}_{1:r}^{i})|<\epsilon/2\quad(n\geq N_{i})$
holds. Let $N$ denote the maximum of $\{N_{i}:i\in[I]\}$. Then, for any $n\geq N$, we have
$\sup_{\bm{a}_{1:r}\in K}|f_{n}(\bm{a}_{1:r})|$ $\leq\max_{i\in[I]}\sup_{\bm{a}_{1:r}\in B_{i}}|f_{n}(\bm{a}_{1:r})|\leq\max_{i\in[I]}\sup_{\bm{a}_{1:r}\in B_{i}}(|f_{n}(\bm{a}_{1:r})-f_{n}(\bm{a}_{1:r}^{i})|+|f_{n}(\bm{a}_{1:r}^{i})|)$
$<\max_{i\in[I]}\epsilon=\epsilon,$
which implies the convergence of $\sup_{\bm{a}_{1:r}\in K}|f_{n}(\bm{a}_{1:r})|$ to zero as $n$ goes to infinity. ∎
C.1.3. $S_{2}$ Converges to 0
We study the convergence of the term $S_{2}$ also defined in Section 4. Specifically, we prove
$S_{2}=\left|\mathbb{E}f\left(\mathbb{E}\left[\psi(Z^{\bm{g}^{1:M}},\bm{p}_{1:r})\mid\bm{p}_{1:r}\right]\right)-\mathbb{E}f\left(\mathbb{E}[\psi(Z^{\bm{g}^{1:M}},\mathring{\bm{p}}_{1:r})\mid\mathring{\bm{p}}_{1:r}]\right)\right|\to 0.$
Define a function $\Psi:\mathbb{R}^{r}\to\mathbb{R}$ by
$\Psi(\bm{a}_{1:r})=\mathbb{E}\left[\psi(Z^{\bm{g}^{1:M}},\bm{a}_{1:r})\right],$
then it is bounded since $\psi$ is bounded. By the bounded convergence theorem, it is also continuous. Therefore, by the continuous mapping theorem, we have
$\Psi(\bm{p}_{1:r})\stackrel{{\scriptstyle d}}{{\longrightarrow}}\Psi(\mathring{\bm{p}}_{1:r}).$
Since $Z^{\bm{g}^{1:M}}$ is independent of $\bm{p}_{1:r}$, we have
$\mathbb{E}\left[\psi(Z^{\bm{g}^{1:M}},\bm{p}_{1:r})\mid\bm{p}_{1:r}\right]=\Psi(\bm{p}_{1:r}).$
Therefore, we haveå
$\mathbb{E}\left[\psi(Z^{\bm{g}^{1:M}},\bm{p}_{1:r})\mid\bm{p}_{1:r}\right]\stackrel{{\scriptstyle d}}{{\longrightarrow}}\Psi(\mathring{\bm{p}}_{1:r}).$
Note that $\Psi(\mathring{\bm{p}}_{1:r})$ can be expressed as
$\Psi(\mathring{\bm{p}}_{1:r})=\mathbb{E}\left[\psi(Z^{\bm{g}^{1:M}},\mathring{\bm{p}}_{1:r})\mid\mathring{\bm{p}}_{1:r}\right],$
where $Z^{\bm{g}^{1:M}}$ is independent of $\mathring{\bm{p}}_{1:r}$. By Lemma 5, this implies $S_{2}\to 0$.
C.2. Proof of Corollary 2
Let $\psi:\mathbb{R}^{J}\to\mathbb{R}$ be a bounded and Lipschitz function that satisfies $|\psi|\leq C$. Note that by Fact 4, it is also pseudo-Lipschitz. Define a bounded and continuous function $f:\mathbb{R}\to\mathbb{R}$ by
$f(x)=-C1\{x<C\}+x1\{x\in[-C,C]\}+C\{x\geq C\}.$
Then, Lemma 5 and Theorem 1 imply that
$\mathbb{E}[\psi(h_{\alpha}^{1},\dots,h_{\alpha}^{J})]=\mathbb{E}\left(\frac{1}{n}\sum_{\alpha=1}^{n}\psi(h_{\alpha}^{1},\dots,h_{\alpha}^{J})\right)=\mathbb{E}\left[f\left(\frac{1}{n}\sum_{\alpha=1}^{n}\psi(h_{\alpha}^{1},\dots,h_{\alpha}^{J})\right)\right]$
$\to\mathbb{E}\left[f\left(\mathbb{E}[\psi(Z^{h^{1}},\dots,Z^{h^{J}})\mid\mathring{p}_{1},\dots,\mathring{p}_{r}]\right)\right]=\mathbb{E}\left[\psi(Z^{h^{1}},\dots,Z^{h^{J}})\right]$
holds as $n\to\infty$. Since the above convergence holds for all bounded and Lipschitz function $\psi$, by Lemma 5, this implies the desired convergence in distribution.